### v1

# Foldable SuperNets: Scalable Merging of Transformers with Different Initializations and Tasks

Abstract
Many recent methods aim to merge neural networks (NNs) with identical architectures trained on different tasks to obtain a single multi-task model. Most existing works tackle the simpler setup of merging NNs initialized from a common pre-trained network, where simple heuristics like weight averaging work well. This work targets a more challenging goal: merging large transformers trained on different tasks from distinct initializations. First, we demonstrate that traditional merging methods fail catastrophically in this setup. To overcome this challenge, we propose Foldable SuperNet Merge (FS-Merge), a method that optimizes a SuperNet to fuse the original models using a feature reconstruction loss. FS-Merge is simple, data-efficient, and capable of merging models of varying widths. We test FS-Merge against existing methods, including knowledge distillation, on MLPs and transformers across various settings, sizes, tasks, and modalities. FS-Merge consistently outperforms them, achieving SOTA results, particularly in limited data scenarios [^1] Code and models are available at https://github.com/idankinderman/fs_merge.
1 Introduction
Practitioners frequently train identical neural architectures for various tasks and share these models online, while the original training data is often unavailable due to privacy, proprietary, or other concerns. This led to an increased interest in the field of model merging (Akhlaghi & Sukhov, 2018; Wortsman et al., 2022; Goddard et al., 2024), which aims to combine the weights, and sometimes the features, of several models into a single new model (Figure 1). This approach could allow the merging of multiple single-task models into a single multi-task model (Matena & Raffel, 2022; Ilharco et al., 2023), eliminating the need to store and run multiple models or ensembles (Ganaie et al., 2022).
Most existing merging methods have a strong restriction: they assume that the models were initialized from the same pre-trained model and subsequently fine-tuned. This encourages the models to stay aligned (Ainsworth et al., 2023) and also to remain closer in the weight space ilharco2023editing, and therefore easier to fuse, for example by simply averaging their weights (wortsman2022model). However, this restricts the ability to merge models that do not share the same initialization. For example, consider the task of merging the weights of two unrelated models from an online repository (e.g., Hugging Face or GitHub). These models were likely not fine-tuned from the same initial model, making most existing merging techniques inapplicable.
To overcome this, a few recent studies explored merging differently initialized models using alignment-based methods (entezari2022the; singh2020model; verma2024merging; ainsworth2023git; stoica2024zipit). However, these approaches use very simple merging rules and hence struggle with larger, complex tasks such as merging transformers (stoica2024zipit).
To illustrate the limitations of traditional and alignment-based merging methods, we pre-trained two Vision Transformers (ViTs) (dosovitskiy2021an) with different initializations on ImageNet-1k (deng2009imagenet), then fine-tuned each on separate tasks: Cars (krause20133d) and CIFAR10 (krizhevsky2009learning). Attempts to merge these models into a single multi-task ViT using various methods (Table 1) were severely unsuccessful. Traditional approaches like weight averaging, SLERP (shoemake1985animating), and RegMean (jin2023dataless), as well as the alignment merging method designed for transformers, ‘Opt’ (imfeld2023transformer), all resulted in a merged model with performance comparable to that of a random guess. Moreover, these methods show a significant accuracy gap compared to the model ensembling (ganaie2022ensemble), a method that averages the model outputs. Note that the ensemble is not a valid merging method, as it uses the original models directly. These traditional merging methods rely on simple local merging rules, ensuring computational efficiency. However, they proved inadequate for our complex setting due to their simplicity. This pattern persisted across all other settings, tasks, and modalities we tested with transformers (see Section 3.2).
Table 1: Merging a pair of ViT-B-16, fine-tuned on Cars and CIFAR10, using 100 original training images and 800 augmented images from each dataset. The test accuracy is averaged on both tasks.
Method Accuracy
Ensemble 89.27
Random guess 5.25
Average 5.56
SLERP 4.80
RegMean 6.58
Opt 6.32
Distillation 75.81
FS-Merge (Ours) 84.52
These results indicate that merging large transformers from diverse initializations demands stronger and more resource-intensive techniques, such as Knowledge Distillation (KD) (ba2014deep; hinton2015distilling). In the multi-task setting, KD refers to a single model learning to replicate the outputs of multiple models (tan2018multilingual; vongkulbhisal2019unifying). And indeed, KD significantly outperforms the previous methods (Table 1), but still exhibits a large performance gap compared to the ensemble. Although KD shows promise, it often requires access to the large parts of the original training dataset and labels (clark2019bam; liu2020adaptive) to achieve high accuracy, which may be problematic due to privacy or proprietary concerns. Moreover, it does not explicitly utilize the original models’ weights, potentially overlooking valuable information.
Figure 1: The model merging setting. (a) Consider models $A$ and $B$ trained from different initializations and on different tasks, which create features processed by a classification head for predictions. (b) Merging methods fuse the models into a new model of the same size while leaving the classification head untouched. The merged model generates features applicable to all tasks.
Our approach. In this work, we address the challenging scenario of merging differently initialized single-task transformer (vaswani2017attention) into a unified multitask model of equal size in a data-efficient manner. We propose Foldable SuperNet Merge (FS-Merge), a method that optimizes a SuperNet to fuse the weights of the original networks. This fusion minimizes local or global feature reconstruction loss. Post-optimization, the weights of the original networks are folded into a single merged model (Figure 2). Importantly, our formulation offers greater generality than simple rule-based alignment techniques (stoica2024zipit; imfeld2023transformer). Moreover, by leveraging the original weights, FS-Merge outperforms KD, particularly in data-limited regimes, significantly reducing the gap with the ensemble (Table 1), and even surpassing it in some cases.
FS-Merge is a simple and data-efficient approach that, like other methods (jin2023dataless; ainsworth2023git), requires only an unlabeled fraction of the original training data. Notably, it produces a merged model that maintains the same inference speed and memory usage as those from traditional methods. We demonstrate its effectiveness by achieving SOTA results across multiple scenarios, architectures, model sizes, datasets, and modalities. While this work focuses on transformers, FS-Merge’s versatility allows it to be readily extended to other architectures such as RNNs (sherstinsky2020fundamentals) and CNNs (he2016deep). This adaptability sets FS-Merge apart from alignment-based methods, which are often designed for specific architectures.
2 Method
Problem formulation.
For simplicity, we outline the merging problem for two models $A$ and $B$ with identical widths, each trained on distinct tasks with unique initializations and separate classification heads (Figure  1a). We also have an unlabeled subset of the training set from each task, $D^{A}$ and $D^{B}$. Our goal is to develop a new model with the same architecture, that minimizes the losses for tasks $A$ and $B$. Note that the classification heads are not merged, meaning we retain a separate classification head for each task (Figure 1b).
In this section, we introduce FS-Merge, our proposed approach. We begin by explaining how to merge MLPs, a relatively easier task, and then proceed to tackle the more complex challenge of merging transformer models.
2.1 Warmup: merging Multi-Layer Perceptrons
Figure 2: The FS-Merge for MLP. (a) Local FS-Merge: $M_{l}$ and $U_{l}$ are optimized separately for each layer $l$ to reconstruct $f_{l}^{A}||f_{l}^{B}$. (b) Global FS-Merge: composing the $M$ and $U$ matrices for each one of the $L$ layers of the MLP to reconstruct $f_{L}^{A}||f_{L}^{B}$. In both versions, after optimization, the Foldable SuperNet is folded to create the merged model. Red represents the SuperNet weights, blue represents the original frozen weights, gray represents features, and green represents the activation function.
We first consider the $l$-th layer in the Multi-Layer Perceptron (MLP) model $A$. The features at this layer, denoted by $f_{l}^{A}\in\mathbb{R}^{d_{l}}$, can be expressed as follows:
$z_{l}^{A}=W_{l}^{A}f_{l-1}^{A}+b_{l}^{A}\,,\,f_{l}^{A}=\sigma(z_{l}^{A})\,.$ (1)
Here, $W_{l}^{A}\in\mathbb{R}^{d_{l}\times d_{l-1}}$ and $b_{l}^{A}\in\mathbb{R}^{d_{l}}$ are the weights and biases of the current linear layer, respectively. $f_{0}^{A}=x$ denotes the MLP input, $d_{l}$ denotes the width of the $l$-th layer, $\sigma$ represents a non-linear element-wise activation function, and $z_{l}^{A}\in\mathbb{R}^{d_{l}}$ are the pre-activation features.
Our method has two versions: local FS-Merge and global FS-Merge. Each has its own Foldable SuperNet and reconstruction optimization problem, neither requiring true labels. The parameters learned during FS-Merge optimization are highlighted in red, and the notation $\tilde{f}$ represents a feature’s reconstruction attempt of the Foldable SuperNet.
Local FS-Merge. In the case of merging the $l$-th linear layers of models $A$ and $B$, the local Foldable SuperNet (Figure 2a) is defined as follows:
$\tilde{f}_{l}(z_{l}^{A},\,z_{l}^{B})={\color[rgb]{1,0,0}U_{l}}\sigma({\color[rgb]{1,0,0}M_{l}}(z_{l}^{A}\,||\,z_{l}^{B}))\,,$ (2)
The input is a concatenation of the original pre-activation features $z_{l}^{A}\,||\,z_{l}^{B}\in\mathbb{R}^{2\cdot d_{l}}$. ${\color[rgb]{1,0,0}M_{l}}\in\mathbb{R}^{d_{l}\times 2\cdot d_{l}}$ (“Merge”) is used to merge the original features into a lower-dimensional space (from $2\cdot d_{l}$ to $d_{l}$), and ${\color[rgb]{1,0,0}U_{l}}\in\mathbb{R}^{2\cdot d_{l}\times d_{l}}$ (“Unmerge”) is used to approximately reverse the merge operation, attempting to reconstruct the original post-activation features ${f}_{l}^{A}\,||\,{f}_{l}^{B}\in\mathbb{R}^{2\cdot d_{l}}$ with $\tilde{f}_{l}\in\mathbb{R}^{2\cdot d_{l}}$. Another way to express $\tilde{f}_{l}$ is by $\tilde{f}_{l}^{A}\,||\,\tilde{f}_{l}^{B}$.
For optimization, $D^{A}\cup D^{B}$ is used to extract the models features $f_{l}^{A},$ and $f_{l}^{B}$. Then, $M_{l}$ and $U_{l}$ are optimized separately for each layer $l$, on the following reconstruction optimization problem:
$M_{l}^{*},U_{l}^{*}=\underset{M_{l},U_{l}}{\text{argmin}}\,\mathbb{E}_{x\sim D}\,\left\|f_{l}^{A}\,||\,f_{l}^{B}-\tilde{f}_{l}(z_{l}^{A},\,z_{l}^{B})\right\|_{2}^{2}\,.$ (3)
Global FS-Merge. In the global version (Figure 2b), the Foldable SuperNet is created by composing the $M$ and $U$ matrices for each one of the $L$ layers of the MLP. In the forward pass, the $l$-th layer of the Foldable SuperNet uses the reconstructed pre-activation features of the previous layer $\tilde{z}_{l}^{A},\,\tilde{z}_{l}^{B}$ as inputs. Observe that this does not include the classification head, which is not being merged. Then, all those matrices are optimized together on the following global optimization problem:
$M_{1}^{*},U_{1}^{*},...M_{L}^{*},U_{L}^{*}=\underset{M_{1},U_{1},...,M_{L},U_{L}}{\text{argmin}}\,\mathbb{E}_{x\sim D}\,\left\|f_{L}^{A}||f_{L}^{B}-\tilde{f}_{L}(x)\right\|_{2}^{2}\,,$ (4)
where $f_{L}^{A}\in\mathbb{R}^{d}$ are features from the last representation layer ($L$-th layer) of model $A$, applied the input $x$. The output of the Foldable SuperNet is defined as $\tilde{f}_{L}\in\mathbb{R}^{2\cdot d}$, which attempts to reconstruct $f_{L}^{A}||f_{L}^{B}\in\mathbb{R}^{2\cdot d_{L}}$. As we will demonstrate later, the global problem allows us to deal with more complicated architectures such as transformers. Note that we have a slight abuse of notation, as $\tilde{f}_{L}$ denotes reconstructed features from both the local and global FS-Merge.
Folding. After optimizing the $M$ and $U$ matrices in all layers (using the local or the global version), we can “fold” the Foldable SuperNet in order to create the merged model. The “folding” operation is defined as follows:
$W_{l}^{*}=M_{l}^{*}\begin{pmatrix}W_{l}^{A}&0\\ 0&W_{l}^{B}\end{pmatrix}U_{l-1}^{*},\ b_{l}^{*}=M_{l}^{*}\begin{pmatrix}b_{l}^{A}\\ b_{l}^{B}\end{pmatrix},\ U_{0}=\begin{pmatrix}I\\ I\end{pmatrix}\,.$ (5)
Intuitively, this folding operation creates a merged model which “under the hood” reconstructs the original features from the previous layer (using $U_{l-1}$), applies the original weights, and then merges those features again ($M_{l}$), all with the same complexity as each of the original models.
Initialization. The Foldable SuperNet can be initialized in various ways (full details in Appendix G.1). In the simplest approach, the $M$ and $U$matrices may be initialized randomly (“random”). We found that the best approach is to initialize so that only the weights of the first model are selected (“first”). This means that, if we fold the weights at that point, we obtain $W^{*}_{l}=W^{A}_{l}$ This is achieved by initializing $M_{l}$ and $U_{l}$ for each layer $l$ as follows:
$M_{l}=\begin{pmatrix}I&0\end{pmatrix}\,,\,U_{l}=\begin{pmatrix}I\\ I\\ \end{pmatrix}.$ (6)
Relation with ZipIt (stoica2024zipit). This “folding” operation was proposed by ZipIt (stoica2024zipit), which is closely related to our work, and inspired it. This method also merges models from various initializations and tasks, targeting MLPs and CNNs. In the MLP context, ZipIt represents a special case of FS-Merge, where $M$ is chosen to average highly correlated pairs in a hard-coded way, and $U=2M^{\top}$. For other layers, such as skip connections and normalizations, ZipIt employs various heuristics that cannot be easily extended to transformers and may harm performance. For more details, see Appendix A.2.
Extension to multiple models. This methodology can easily be extended to any number of models. It is also capable of merging models of varying widths into any target width dimension, provided they have the same number of layers.
2.2 Merging transformers
Figure 3: The Foldable SuperNet for merging two attention blocks ($A$ and $B$). Only the red components are trained. After training, we fold the Foldable SuperNet to create the merged block.
Merging transformers (vaswani2017attention) is much more challenging than merging MLPs, due to their larger scale and more complicated structure. Creating a naively Foldable SuperNet as described in the MLP Section 2.1 is not feasible for transformers because it fails to account for their skip connections, layer normalization (ba2016layer), and multi-head attention.
To tackle these issues, we develop a new Foldable SuperNet architecture, and train it using the global objective Eq. 4 (see Appendix B.2 for more details). We found that the global objective is essential in the case of transformers, as training the Foldable SuperNet locally (Eq. 3) significantly reduces the accuracy of the resulting merged model (in Appendix H.1 we suggest a few reasons for this). In this section, we explain how to build a Foldable SuperNet for the attention block (Figure  3), as it is the most complex component of the transformer. For the pre-processing block and the MLP block, please refer to Appendix B.1.
A review of the attention block. Consider the transformer trained on task $A$. The $l$-th attention block gets the features $X_{l}^{A}\in\mathbb{R}^{T\times d}$ as input, when $T$ is the sequence length and $d$ is the embedding size, and applies layer norm with parameters $\gamma^{A}_{l},\beta^{A}_{l}\in\mathbb{R}^{d}$ (Eq. 7 left). Then, it calculates the queries, keys, and values for $H$ heads. An example of query creation is shown in Eq. 7 right.
$\bar{X}_{l}^{A}=\text{LN}_{\gamma^{A}_{l},\beta^{B}_{l}}(X_{l}^{A})\,\,\,;\,\,\,Q^{A}_{l}=\bar{X}_{l}^{A}W_{l}^{Q,A}=(Q_{1,l}^{A},...,Q_{H,l}^{A})\,.$ (7)
Where $Q_{i,l}^{A}\in\mathbb{R}^{T\times\frac{d}{H}}$ is the $i$-th head queries. Similarly, the keys and values are created using $W_{l}^{K,A},\,W_{l}^{V,A}$ respectively. Following this, the transformer executes multi-head attention, utilizes $W_{l}^{O,A}$, and uses a skip connection, which creates the output $Y_{l}^{A}\in\mathbb{R}^{T\times d}$
$Y_{l}^{A}=X_{l}^{A}+\text{Concat}_{i}[...,\text{softmax}\left(\frac{Q_{i,l}^{A}{K_{i,l}^{A}}^{\top}}{\sqrt{d}}\right)V_{i,l}^{A},...]W_{l}^{O,A}\,.$ (8)
Foldable SuperNet for attention blocks. Suppose there are two transformers trained on distinct tasks $A$ and $B$. Our goal is to define a Foldable SuperNet that will use the original weights and will allow us to merge the attention blocks of these two transformers after training. Intuitively, this Foldable SuperNet will reconstruct the original features of models $A$ and $B$, create queries keys and values, learn how to merge them, and apply multi-head attention. The optimized parameters at this stage are highlighted in red. The notation $\tilde{X}$ represents a reconstruction attempt of our method, and $X^{*}$ represents the merged features or parameters.
The $l$-th attention block gets merged features from the previous layer as input, $X_{l}^{*}\in\mathbb{R}^{T\times d}$. Then, the Foldable SuperNet learns the parameters of the layer norm of this block ${\color[rgb]{1,0,0}\gamma^{*}_{l},\beta^{*}_{l}}\in\mathbb{R}^{d}$, along with a matrix ${\color[rgb]{1,0,0}U_{l}}\in\mathbb{R}^{d\times 2\cdot d}$ to reconstruct the features from the merged input
$\tilde{X}_{l}^{A}||\tilde{X}_{l}^{B}=\text{LN}_{{\color[rgb]{1,0,0}\gamma^{*}_{l},\beta^{*}_{l}}}(X_{l}^{*}){\color[rgb]{1,0,0}U_{l}}\,.$ (9)
Observe that the layer norm in each block is also being optimized, similar to what has been proposed in other merging works (jordan2023repair). This does not hurt the efficiency of our method due to the small number of learnable parameters in the layer norm. Then, the Foldable SuperNet calculates the queries, keys, and values for $A$ and $B$. These components, drawn from all heads and models, are then concatenated and merged using learnable matrices ${\color[rgb]{1,0,0}M_{l}^{Q},M_{l}^{K},M_{l}^{V}}\in\mathbb{R}^{2\cdot d\times d}$. For example, the queries are merged as follows:
$\tilde{Q}^{A}_{l}||\tilde{Q}^{B}_{l}=(\tilde{X_{l}}^{A}W_{l}^{Q,A}||\tilde{X_{l}}^{B}W_{l}^{Q,B})\in\mathbb{R}^{T\times 2\cdot d}\,,\,Q_{l}^{*}=(\tilde{Q}^{A}_{l}||\tilde{Q}^{B}_{l}){\color[rgb]{1,0,0}M_{l}^{Q}}=(Q_{1,l}^{*},...,Q_{H,l}^{*})\,.$ (10)
Then, they are divided into $H$ heads. Similarly, the keys and values are compressed using the matrices ${\color[rgb]{1,0,0}M_{l}^{K},M_{l}^{V}}$ respectively. Following this, the Foldable SuperNet executes multi-head attention in the original width of $d$, thanks to the compression of queries, keys and values; uses ${\color[rgb]{1,0,0}U_{l}^{O}}\in\mathbb{R}^{d\times 2\cdot d}$ to reconstruct the original multi-head attention output; utilizes $W_{l}^{O,A}\,,\,W_{l}^{O,B}$; uses ${\color[rgb]{1,0,0}M_{l}^{O}}\in\mathbb{R}^{2\cdot d\times d}$ to compress it once more; and applies a skip connection. Together, we get:
$Y_{l}^{*}=X_{l}^{*}+\text{Concat}_{i}[...,\text{softmax}\left(\frac{Q_{i,l}^{*}{K_{i,l}^{*}}^{\top}}{\sqrt{d}}\right)V_{i,l}^{*},...]\ {\color[rgb]{1,0,0}U_{l}^{O}}\begin{pmatrix}W_{l}^{O,A}&0\\ 0&W_{l}^{O,B}\end{pmatrix}{\color[rgb]{1,0,0}M_{l}^{O}}\,.$ (11)
This process results in $Y_{l}^{*}\in\mathbb{R}^{T\times d}$, serving as the input for the subsequent MLP block at layer $l$. It is important to note that this Foldable SuperNet was designed so the skip connection and layer norm are applied to the compressed features, and that for any $M$ and $U$, there is a linear layer that comes before or after it. This arrangement will allow us to fold this structure into the merged model after training (see Appendix B.3 for more details).
Parameterizing $M$ and $U$. Utilizing full-rank $M_{l}\in\mathbb{R}^{d_{l}\times n\cdot d_{l}}$ and $U_{l}\in\mathbb{R}^{n\cdot d_{l}\times d_{l}}$ for merging $n$ models introduces a number of parameters that increase quadratically with the layer width $d_{l}$. In large models like transformers, this leads to a very high demand for resources and hinder the optimization process. To mitigate this, we adopt a parameterization strategy akin to LoRA (hu2022lora), using a sum of a low-rank matrix and a concatenation of diagonal matrices. For instance, in $M_{l}$:
$M_{l}=M_{l}^{\text{diag}}+M_{l}^{1}M_{l}^{2}\,.$ (12)
When $r$ is the inner rank, $M_{l}^{1}\in\mathbb{R}^{d_{l}\times r}$, $M_{l}^{2}\in\mathbb{R}^{r\times n\cdot d_{l}}$, and $M_{l}^{\text{diag}}\in\mathbb{R}^{d_{l}\times n\cdot d_{l}}$ is a concatenation of a $n$ diagonal matrices, each with $d_{l}$ learnable parameters. A similar structure is proposed for $U_{l}$. This ensures the number of learnable parameters is linear with the layer width $d_{l}$. Additionally, we found that adding $M_{l}^{\text{diag}}$ is crucial for FS-Merge, as it enables initializing the Foldable SuperNet with strong initializations such as “first” (Eq. 6).
FS-Merge seq. To address the high costs of merging a large number of models, we introduce a more efficient variant called FS-Merge Seq., which merges the models sequentially. It starts with the first two models, then continues by merging the resulting merged model with the third model, and so on. Full details can be found in Appendix B.4.
Data and Augmentation. This work addresses a realistic and challenging setting, involving a limited subset of unlabeled samples used for merging. For transformer merges, augmentations (zhang2018mixup) are employed to expand this subset, as commonly done in regular training. Appendix G.1 studies the effect of using augmentation on accuracy.
3 Results
We evaluate our method on MLPs (Section 3.1), Vision Transformers (Section 3.2), and Text Transformers (Appendix F.3), and show it achieves state-of-the-art results.
Baselines. We compared with “Original Models”, representing the average accuracy of the models to be merged; and Ensemble (ganaie2022ensemble), which averages the models outputs and then applies classification heads. Note that these are not valid merging methods as they use the original models directly. For legitimate merging techniques, comparisons were made with weight averaging (“average”) (wortsman2022model); RegMean (jin2023dataless) which applies a closed-form linear regression solution to each layer; and distillation (hinton2015distilling) which trains a single model to mimic the pre-classification layer features. ZipIt (stoica2024zipit) averages highly correlated neurons, used only in the MLP experiments, as it is inapplicable for transformers. “Opt” (imfeld2023transformer), uses optimal transport (knight2008sinkhorn) for aligning transformers, and “SLERP” (spherical linear interpolation) (shoemake1985animating), utilized only on the ViT case, as they introduced for transformers.
Metrics. As in ZipIt (stoica2024zipit), we evaluate multi-task merged models using two metrics: per-task accuracy and joint accuracy. In per-task accuracy, we calculate the accuracy for each task individually using only the relevant classification head. Then the mean accuracy across all tasks is reported. In joint accuracy, we calculate the accuracy for each task by making predictions based on the maximum score across all classification heads and reporting the mean accuracy across the tasks.
3.1 Merging Multi-Layer Perceptrons
Table 2: We merged pairs of MLPs, each initialized differently and trained on distinct halves of the MNIST dataset. These MLPs have a hidden width of 128 neurons, with the number of hidden layers varying from 1 to 6. Each experiment was replicated five times. We present the average per-task accuracy on the test set, along with the standard deviation.
Merge Method Number of Hidden Layers
1 2 3 4 6
Original Models 96.83 $\pm$ 0.13 96.51 $\pm$ 0.19 96.46 $\pm$ 0.23 95.8 $\pm$ 0.4 96.8 $\pm$ 0.5
Ensemble 94.70 $\pm$ 0.95 95.14 $\pm$ 1.12 95.73 $\pm$ 0.22 95.5 $\pm$ 0.1 95.7 $\pm$ 0.7
Average 94.36 $\pm$ 0.76 85.90 $\pm$ 4.46 78.78 $\pm$ 6.72 61.7 $\pm$ 6.3 25.1 $\pm$ 2.7
RegMean 95.90 $\pm$ 0.37 92.97 $\pm$ 2.71 92.11 $\pm$ 1.90 87.7 $\pm$ 3.7 81.5 $\pm$ 2.4
ZipIt 96.35 $\pm$ 0.17 95.75 $\pm$ 0.58 95.43 $\pm$ 0.50 94.5 $\pm$ 0.5 94.0 $\pm$ 2.2
Distillation 93.35 $\pm$ 1.07 93.13 $\pm$ 1.74 93.71 $\pm$ 0.39 93.3 $\pm$ 0.6 90.7 $\pm$ 1.8
FS-M 95.89 $\pm$ 0.03 95.68 $\pm$ 0.19 95.37 $\pm$ 0.33 94.9 $\pm$ 0.4 94.8 $\pm$ 0.8
FS-M, ZipIt init 96.62 $\pm$ 0.08 96.29 $\pm$ 0.24 96.18 $\pm$ 0.20 95.6 $\pm$ 0.5 96.2 $\pm$ 0.8
Setting. Our evaluation started with a straightforward experiment merging pairs of MLPs. The MLPs were trained on divided MNIST (lecun1998mnist), meaning it was split into two subsets: images with labels 0-4 and images with labels 5-9. These subsets were further divided into training, validation (10% of the training set), and test sets. Two MLPs were trained separately on these subsets, varying in the number of layers and widths, each initialized with a different seed. The goal is to merge these pairs of models. Note that we did not merge the last linear layer, which acts as a classification head.
FS-Merge. Our method was tested in two variants: FS-Merge in the local version (Eq. 3), which trains a Foldable SuperNet for each layer independently; and FS-Merge ZipIt, which initializes the Foldable SuperNet’s $M$ and $U$ as the solutions of ZipIt (stoica2024zipit), and then optimize them using the local version. Data. All merging methods, excluding “Average”, use features from the original models. Thus, we sampled 64 images from each dataset’s training set to generate the necessary features.
Table 2 presents the per-task accuracy on the test set, for merging MLPs trained on half of the MNIST dataset. We merged MLPs with 128 hidden widths, and hidden layers varying from 1 to 6. Each experimental condition was replicated five times with 5 different seeds. The same hyperparameters were used for all those experiments. Full information about the setting and hyperparameters are available in Appendix E.1.
Our results indicate that merging deeper models is more challenging, consistent with previous studies (jordan2023repair). Employing the ZipIt initialization, our method establishes a new SOTA for both per-task and joint accuracy, outperforming ensemble in many cases, and nearly matches the accuracy of “Original Models”. For results on more tasks and FS-Merge versions, see Appendix F.1
Table 3: Merging pairs of ViT-B-16 using 16 original images from each training set and 800 augmented images from each dataset. The per-task and joint accuracy on the test set are reported.
Merging Methods DTD, EuroSAT CIFAR100, SVHN RESISC45, SVHN
Per-task Joint Per-task Joint Per-task Joint #Parameters Optimized
Original models 81.55 - 91.19 - 95.12 - -
Ensemble 78.64 74.11 88.26 57.44 93.21 75.07 -
Average 11.48 1.15 3.71 0.84 6.23 3.68 0
SLERP 7.99 1.27 4.10 0.65 6.67 3.24 0
RegMean 8.40 1.69 6.17 1.27 6.44 0.74 0
Opt 4.51 0.75 4.52 0.88 7.01 2.07 0
Distillation 57.31 52.61 62.93 48.12 66.18 63.16  [^1] M
FS-M 63.18 59.28 66.02 49.42 73.43 69.11 13M
3.2 Merging Vision Transformers
Next, we evaluated our method on merging Vision Transformers (ViT), which were initialized differently and trained on distinct tasks— a much more challenging setting. Models and Data. We pre-trained several ViT-B-16 and ViT-L-14 models (dosovitskiy2021an; touvron2021training) on ImageNet-1K (deng2009imagenet). These models were initialized from distinct random seeds and exposed to training data in varying sequences. Then, each differently pre-trained ViT was fine-tuned on downstream tasks. Following (ilharco2023editing), we fine-tuned on Cars (krause20133d), DTD (cimpoi2014describing), EuroSAT (helber2019eurosat), GTSRB (stallkamp2011german), MNIST (lecun1998mnist), RESISC45 (cheng2017remote), SVHN (netzer2011reading), CIFAR10, and CIFAR100 (krizhevsky2009learning). For extended details regarding the pre-training, datasets and finetuning please refer to Appendix D.
In these experiments, the KD baseline also used the “first” initialization, meaning that the student model was initialized from the first model. We found it to outperform any other initialization, including using traditional merging methods such as “RegMean” for initialization and then applying distillation. See Appendix G.2 for full details.
FS-Merge. We use the global version of FS-Merge, when the whole Foldable SuperNet is trained to reconstruct the features of the original models from the layer preceding the classification head. The Foldable SuperNet’s $M$ and $U$ were parametrized as a concatenation of diagonal matrices plus low-rank matrices (Eq. 12). As in the distillation method, the Foldable SuperNet was initialized using the “first” initialization (Eq. 6), as we found it has the best performance (Appendix G.1). FS-M Seq. (Appendix B.4) is a memory and compute-efficient version of FS-Merge, specifically designed for merging a large number of models.
Table 4: Merging groups of 4 ViT-B-16 with 100 original images from the training set and 1000 augmented images from each dataset (a total of 400 original images and 4,000 augmented images). We report the per-task and joint accuracy on the test set. We will denote: C = Cars, D = DTD, E = EuroSAT, G = GTSRB, M = MNIST, R = RESISC45, S = SVHN, C10 = CIFAR10, C100 = CIFAR100.
Merging Methods R, C10, S, G D, G, E, R C, M, C100, E
Per-task Joint Per-task Joint Per-task Joint #Parameters Optimized
Original models 96.47 - 88.72 - 92.61 - -
Ensemble 86.11 46.80 76.81 52.96 82.04 63.92 -
Average 5.40 1.04 3.66 1.35 4.55 0.38 0
SLERP 5.55 1.04 4.88 0.87 7.27 0.86 0
RegMean 6.38 0.61 4.78 0.61 5.60 0.52 0
Opt 5.79 0.24 3.70 0.38 5.75 2.58 0
Distillation 82.09 67.60 67.31 57.67 37.83 31.71  [^1] M
FS-M 84.34 71.17 67.43 55.35 79.48 71.24 60M
FS-M seq. 83.11 70.34 67.55 56.90 73.13 66.03 18M
In Table 3, pairs of ViT-B-16 models fine-tuned on different tasks were merged, in a low-data scenario of using only 16 original images per dataset. An additional 800 augmented images per dataset were created. FS-merge was used with a low rank of 12. To examine the effect of merging a larger number of models, Table 4 shows the results for merging groups of four ViT-B-16 models. FS-merge was used with a low rank of 32, and FS-Merge seq. was used with a low rank of 16. To evaluate the impact of merging larger models, Table 5 presents the results of merging pairs of ViT-L-14 models. FS-merge was used with a low rank of 32. For extended ViT merging results in all these settings and more, and hyperparameter specifics, see Appendix F.2 and Appendix E.2.
Table 5: Merging pairs of ViT-L-14 with 100 original images from the training set and 1000 augmented images from each dataset. We report the per-task and joint accuracy on the test set.
Merging Methods DTD, EuroSAT CIFAR100, SVHN Cars, MNIST
Per-task Joint Per-task Joint Per-task Joint #Parameters Optimized
Original models 81.33 - 94.68 - 96.50 - -
Ensemble 77.71 71.54 94.25 77.90 96.23 96.22 -
Average 6.36 1.40 9.18 8.03 5.98 0.08 0
SLERP 5.21 1.58 5.20 2.59 8.48 5.44 0
RegMean 8.66 4.21 5.64 0.47 10.97 0.13 0
Opt 10.33 3.11 4.68 2.45 5.95 5.67 0
Distillation 78.51 75.84 90.91 85.78 91.82 90.58 342M
FS-M 78.60 74.86 91.68 90.92 95.77 95.09 95M
Discussion. As can be seen, FS-Merge outperforming all other merging methods in most cases, and even surpassing ensembles in some cases. This holds for both per-task and joint accuracy across all settings, despite using fewer learnable parameters than distillation. Additionally, it is evident that all local and simple methods (such as Average, SLERP, RegMean, and “Opt”) completely fail to effectively merge ViTs in this challenging setting, resulting in a merged model that performs comparably to a random guess. We also find that FS-Merge achieves SOTA results when merging BERTs on NLP tasks, as can be seen in Appendix F.3.
Our experiments indicate that in the ViT case, initialization is crucial for FS-Merge as it does not converge when initialized randomly (and see Appendix H.2). Specifically, using the “first” initialization in Foldable SuperNet, and KD, not only improves the accuracy of the first task but also enhances accuracy across all tasks. For an ablation study in the matter, see Appendix G.1 and Appendix G.2.
Merging Complexity. FS-Merge and KD are more computationally intensive than standard merging methods such as Averaging and RegMean, which fail catastrophically in our setting. Notably, the accuracy of both FS-Merge and KD cannot be improved simply by longer training (i.e. more resources), as this causes overfitting in our data-scarce regime (this explains why the longest training duration was not identified as optimal in our hyperparameter search). When merging two models, FS-Merge and distillation have comparable resource usage. However, when merging multiple models, the resource usage gap between FS-Merge and distillation becomes more significant. To address this, we propose FS-Merge seq., which is comparable to distillation’s resource use while outperforming it in terms of accuracy. Therefore, FS-Merge is recommended for optimal test accuracy, given sufficient computational resources and limited data. If resources are more limited, FS-Merge seq. should be the method of choice. Refer to Appendix C for the full details.
Figure 4: We used Ensemble, Distillation, and FS-Merge to merge pairs of models trained on RESISC45 and GTSRB (left), Cars and MNIST (center), DTD and CIFAR10 (right). We varied the number of original images per dataset and added augmentation images so the total number of images per dataset would be 1024. We present the per-task and joint accuracy.
3.3 Number of original training images
We examine the impact of varying $|D|$, the number of images taken from the training datasets of the models to be merged (“original images”), used to create features. We varied $|D|$ from 16 to 1024. Augmented images were created to ensure the total number of images per dataset reached 1024, thus maintaining a consistent dataset size. Pairs of ViT-B-16 models were merged using Ensemble, Distillation, and FS-Merge with low rank of 24. The per-task and joint accuracies on the test set are presented in Figure 4. For additional details and experiments, refer to Appendix F.4.
As observed, Distillation underperforms with few original images, while FS-Merge excels. Increasing original images enhances all techniques, reducing the performance gap. With enough data, merging methods can sometimes surpass ensemble performance, which is often considered as a “gold-standard method” in merging and multitask articles.
We report that FS-Merge and Distillation achieve perfect fit on the original images in all cases. We argue that low-rank FS-Merge attains better generalization than distillation due to a useful inductive bias, which constrains the merged model to be a low-rank weighted average of the original model’s weights (as determined by the Foldable SuperNet).
4 Conclusion
Limitations. FS-Merge requires a small unlabeled subset of the original training data, similarly to most previous merging methods. Additionally, FS-Merge, like distillation, is more computationally intensive than standard merging methods; however, these methods completely failed in more challenging settings. Notably, the main bottleneck in model merging is often data availability rather than computational resources. Moreover, FS-Merge has fewer learnable parameters than distillation, but they increase linearly with the number of models and hidden width; however, FS-Merge seq. solved the first issue. Lastly, one cannot naively merge two models of different depths using our method; we believe this could be solved in future work.
Summary. In this work, we address the challenging task of merging transformers from different initializations and tasks into a unified multitask model using a small subset of unlabeled data—a setting in which traditional methods fail. Our proposed FS-Merge, which employs a feature reconstruction approach to train a Foldable SuperNet, is simple, data-efficient, and can use more sophisticated merging rules compared to other baselines. FS-Merge outperforms traditional methods and achieves SOTA results across various scenarios, model sizes, datasets and modalities.
Reproducibility. The paper fully discloses all the information needed to reproduce the experimental results. The method is detailed in Section 2; the main results are shown in Section 3; the pre-training details are explained in Appendix D; the full experimental details and hyperparameters are written in Appendix E; and the additional results can be found in Appendix F.1. In addition, code will be published upon acceptance.
Ethics. We can identify several aspects where FS-Merge can mitigate ethical concerns. The ability to merge models from different initializations and tasks offers an efficient alternative for using an ensemble of these models. This allows us to achieve greater resource efficiency and reduce the model’s carbon footprint. Moreover, FS-Merge can alleviate privacy concerns. For example, in cases where multiple users are training models on private datasets, FS-Merge enables us to combine those models into a single multi-task model, without accessing the full private datasets or any labels.
Acknowledgements
The research of DS was Funded by the European Union (ERC, A-B-C-Deep, 101039436). Views and opinions expressed are however those of the author only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency (ERCEA). Neither the European Union nor the granting authority can be held responsible for them. DS also acknowledges the support of the Schmidt Career Advancement Chair in AI. HM is the Robert J. Shillman Fellow, and is supported by the Israel Science Foundation through a personal grant (ISF 264/23) and an equipment grant (ISF 532/23).
Appendix A Related work
A.1 Related work
Mode Connectivity.
A pair of models has mode connectivity when there exists a simple path between them in the loss landscape with low loss (garipov2018loss; draxler2018essentially; lubana2023mechanistic). frankle2020linear demonstrated that this path can be linear (LMC) when the models originate from the same initialization. tatro2020optimizing; ainsworth2023git; entezari2022the showed that it is feasible to establish LMC between MLPs and CNNs from different initializations using permutations. jordan2023repair enhanced these solutions by addressing their variance collapse. Wang2020Federated used these permutations to match and then merge models within a federated learning framework. imfeld2023transformer; verma2024merging used permutations to align two transformers with different initializations that were trained on the same task. mirzadeh2021linear leveraged LMC for multitask and continual learning. yunis2022on expanded LMC to more than two models, discovering a high-dimensional convex hull of low loss. zhou2023going introduced Layerwise Linear Feature Connectivity (LLFC), demanding that the features of the models be linearly connected. singh2020model; akash2022wasserstein addressed a similar challenge of neuron alignment using optimal transport (knight2008sinkhorn), and liu2022deep generalized it as a graph-matching task. navon2023equivariant solved the neuron alignments problem by training an equivariant deep weight space network. It is important to note that previous works which merged models from different initializations using permutations primarily focused on merging pairs of models trained on the same task, with many of them concentrating on MLPs and CNNs. In contrast, our method is capable of merging larger groups of models, with different initializations and tasks, and is specifically designed to handle transformers.
Model merging.
Model merging technique (goddard2024arcee) has gained increasing interest in the past years, allowing the creation of stronger single-task and multi-task models. akhlaghi2018knowledge showed that averaging the weights of multiple simple neural networks maintains their performance. Model soups (wortsman2022model) proposed averaging multiple models trained on the same task from identical initializations to enhance task accuracy. In addition, Weight averaging has been employed for various purposes, including improving optimization through checkpoint averaging (izmailov2018averaging), federated learning (mcmahan2017communication), developing superior pre-trained models (choshen2022fusing; don2022cold), improving Out-of-Distribution Generalization (rame2022diverse), achieving success in unseen tasks (huang2023lorahub), and as augmentations for weight space networks (shamsian2024improved). In many cases, spherical linear interpolation (SLERP) (shoemake1985animating) is used to average the models’ weights. matena2022merging utilized Fisher-Weighted Averaging to fuse multiple models from the same initializations but trained on diverse tasks, resulting in a multi-task model. RegMean (jin2023dataless) addressed a similar scenario and proposed a closed-form solution that solves a local linear regression problem for each linear layer in the model. ilharco2023editing defined task vector by subtracting the parameters of a fine-tuned model from those of the pre-trained model, and used it to fuse models fine-tuned from the same pre-trained model. yadav2024ties; ortiz2024task; yang2024adamerging; akiba2024evolutionary analyzed and proposed more merging methods based on task vectos. sung2023an fused pairs of models trained on different modalities. ZipIt (stoica2024zipit) merged models from various initializations and tasks, focusing on MLPs and CNNs, by averaging pairs of highly correlated neurons between and within the models (see Appendix A.2 for more details). In contrast, our work focuses on MLPs and transformers, and can use much more complicated merging schemes.
Distillation.
In Knowledge Distillation (KD), a small student model is trained to mimic the outputs of a larger teacher model (ba2014deep; hinton2015distilling; stanton2021does; gou2021knowledge). sau2016deep implemented a noise-based methodology to simulate learning a single task from multiple teachers. tan2018multilingual; khanuja2021mergedistill used several teachers, each translating between a specific language pair, to train a singular multilingual student. clark2019bam; park2023multitask; chelaramani2021multi utilized multiple models and true labels to instruct a multi-task model. pham2023collaborative employed numerous quantized teachers trained on the same task to teach a quantized model, also leveraging true labels. wu2021one co-fine-tuned multiple teachers in downstream tasks with shared layers to instruct a student. jacob2023online simultaneously trained multiple single-task models with a single multi-task student to facilitate the student’s optimization. Similar to our distillation baseline, vongkulbhisal2019unifying utilized KD to unify knowledge from various models with different targets into a single model, relying solely on unlabeled data. wu2021one; zagoruyko2017paying; heo2019knowledge; heo2019comprehensive focused on using a single teacher’s outputs and inner features to train a single model. park2019feed; liu2020adaptive applied multiple teachers trained on the same task, their inner features, and the true labels to train a single-task student. Inspired these works, we tried to merge transformers using a loss function that takes into account the inner features. However, this approach proved ineffective for both our method and distillation. See Appendix H.3 for details.
A.2 Comparison to ZipIt
Our work strongly relates to ZipIt (stoica2024zipit), which also merges models from various initializations and tasks, with a focus on MLPs and CNNs. We will briefly explain the algorithm and its limitations. ZipIt uses the original training data to extract features from the models to be merged, $A$ and $B$. Then, for each layer $l$, it employs the features to identify pairs of highly correlated neurons using a greedy algorithm. Following this, it builds $M_{l}\in\mathbb{R}^{d_{l}\times 2\cdot d_{l}}$, which is zero except for entries corresponding to a matched pair $(i,j)$ indexed by $p$, where $M_{l}[p,i]=M_{l}[p,j]=\frac{1}{2}$. The purpose of this matrix is to merge features by averaging pairs of highly correlated neurons. Then, $U_{l}=2M_{l}^{\top}$ is used to reposition the merged features back to their original locations. After building $M_{l}$ and $U_{l}$, ZipIt proposes folding them with the original weights (Eq. 5) to create the weights of the merged model.
Like other current model merging methods (ainsworth2023git; jin2023dataless; imfeld2023transformer), ZipIt is efficient and focuses only on simple fusion schemes. It limit itself to pairing similar neurons, addressing only the permutation symmetries of neural networks (hecht1990algebraic). Permutation symmetries mean that it is possible to swap any two neurons of a hidden layer in a neural network without altering its functionality. However, ZipIt falls short in handling the scale symmetries of neural networks (neyshabur2015path; badrinarayanan2015understanding; phuong2019functional), or in considering more complicated merging rules.
Probably due to these limitations, ZipIt underperforms on a large scale (such as ResNet-50 trained on datasets with 200 categories). Furthermore, this merging method is formulated as local problems, merging one layer at a time, and relies on heuristics to handle more complicated layers (such as batch normalization and skip connections). This makes it difficult to generalize to more complex architectures like transformers, where ZipIt struggles with self-attention and skip connection structures.
Our work aims to solve a similar problem, but adopts a more expressive approach without relying on heuristics. Moreover, the global version of our method allows the merging of any architecture by simply constructing a Foldable SuperNet that is suitable for it.
Appendix B Merging Vision Transformers with FS-Merge
B.1 Foldable SuperNet for Vision Transformers
Assuming there are two Vision Transformers (ViTs) (dosovitskiy2021an), trained on two distinct tasks, $A$ and $B$, we aim to define a Foldable SuperNet that combines the original weights of the ViTs with new learnable parameters $M,U$. This structure is designed so that all skip connections and layer normalizations (ba2016layer; xiong2020layer) operate on the merged dimension, and that there is a linear layer before or after every $M$ or $U$ matrix, allowing them to be folded after training. It is important to highlight that layer norms possess significantly fewer learnable parameters compared to other layers in ViTs. Therefore, we can initiate their parameters from a good starting point (e.g., the parameters of the first model to be merged) and proceed to optimize the parameters as usual, similar to strategies employed in previous merging works (jordan2023repair; stoica2024zipit).
The method described here for merging two models can be readily extended to any number of models. The parameters optimized at this stage are highlighted in red. The notation $\bar{X}$ represents the outputs of a layer norm, $\tilde{X}$ represents a feature reconstruction attempt after using the $U$ matrix, and $X^{*}$ represent the merged features.
B.1.1 Pre-processing
First, each ViT creates patches from the input image and reshapes them into a series of vectors $I_{\text{proj}}\in\mathbb{R}^{T-1\times d_{\text{in}}}$, using $W^{\text{in}}\in\mathbb{R}^{d_{\text{in}}\times d}$ to project these vectors into tokens. It then concatenates the $\text{CLS}\in\mathbb{R}^{d}$ token and adds $\text{emb}\in\mathbb{R}^{T\times d}$ which is the positional encoding. For example, for model $A$:
$Z^{A}=\text{Concat}[I_{\text{proj}}W^{\text{in},A}\,,\,\text{CLS}^{A}]+\text{emb}^{A}\,.$
The tokens from both models are then concatenated to form $(Z^{A}||Z^{B})\in\mathbb{R}^{T\times 2\cdot d}$, and the Foldable SuperNet merges them using a learned matrix ${\color[rgb]{1,0,0}M_{\text{in}}}\in\mathbb{R}^{2\cdot d\times d}$ to produce $Z^{*}\in\mathbb{R}^{T\times d}$.
$Z^{*}=(Z^{A}||Z^{B}){\color[rgb]{1,0,0}M_{\text{in}}}\,.$
Subsequently, and like the ViT, the Foldable SuperNet applies layer normalization, with parameters ${\color[rgb]{1,0,0}\gamma^{*}_{-1},\beta^{*}_{-1}}\in\mathbb{R}^{d}$, to generate the input for the transformer. These parameters will be optimized.
$\bar{X}_{0}^{*}=\text{LN}_{{\color[rgb]{1,0,0}\gamma^{*}_{-1},\beta^{*}_{-1}}}(Z^{*})$
B.1.2 The attention block
The Foldable SuperNet of the attention block at layer $l$ (Figure  3) receives merged features $X^{*}_{l}\in\mathbb{R}^{T\times d}$ as inputs, and applies layer normalization. It’s parameters ${\color[rgb]{1,0,0}\gamma^{*}_{l},\beta^{*}_{l}}\in\mathbb{R}^{d}$ will be learned. Additionally, a learnable matrix ${\color[rgb]{1,0,0}U_{l}}\in\mathbb{R}^{d\times 2\cdot d}$ is introduced to reconstruct the original features from the merged ones.
$\tilde{X}_{l}^{A}||\tilde{X}_{l}^{B}=\text{LN}_{{\color[rgb]{1,0,0}\gamma^{*}_{l},\beta^{*}_{l}}}(X_{l}^{*}){\color[rgb]{1,0,0}U_{l}}\,.$
After the layer normalization, the queries, keys, and values of each ViT are calculated. Then the Foldable SuperNet concatenates these components, from all heads and models, and merges them using a learnable matrix. Taking the queries as an example:
$\tilde{Q}^{A}_{l}||\tilde{Q}^{B}_{l}=(\tilde{X_{l}}^{A}W_{l}^{Q,A}||\tilde{X_{l}}^{B}W_{l}^{Q,B})\in\mathbb{R}^{T\times 2\cdot d}\,,$
$Q_{l}^{*}=(\tilde{Q}^{A}_{l}||\tilde{Q}^{B}_{l}){\color[rgb]{1,0,0}M_{l}^{Q}}=(Q_{1,l}^{*},...,Q_{H,l}^{*})\,.$
The weights $W_{l}^{Q,A},W_{l}^{Q,B}\in\mathbb{R}^{d\times d}$ generate queries from the embeddings. The matrix ${\color[rgb]{1,0,0}M_{l}^{Q}}\in\mathbb{R}^{2\cdot d\times d}$ merges these features, which are then divided into $H$ heads, where each head $Q_{1,l}^{*}$ has dimensions $\mathbb{R}^{T\times\frac{d}{H}}$. Similarly, the keys and values are created by $W_{l}^{K},W_{l}^{V}\in\mathbb{R}^{d\times d}$, and compressed using the matrices ${\color[rgb]{1,0,0}M_{l}^{K},M_{l}^{V}}\in\mathbb{R}^{2\cdot d\times d}$ respectively.
Following this, and similar to the ViT, the Foldable SuperNet executes multi-head attention with the merged queries, keys and values and concatenates the features from the heads. In our Foldable SuperNet, this step also includes adding ${\color[rgb]{1,0,0}U_{l}^{O}}\in\mathbb{R}^{d\times 2\cdot d}$ to reconstruct the original multi-head attention outputs. Then each ViT utilizes $W_{l}^{O}\in\mathbb{R}^{d\times d}$ to aggregate those outputs. We also use ${\color[rgb]{1,0,0}M_{l}^{O}}\in\mathbb{R}^{2\cdot d\times d}$ to compress it once more. This is followed by a skip connection.
$Y_{l}^{*}=X_{l}^{*}+\text{Concat}_{i}[...,\text{softmax}\left(\frac{Q_{i,l}^{*}{K_{i,l}^{*}}^{\top}}{\sqrt{d}}\right)V_{i,l}^{*},...]\ {\color[rgb]{1,0,0}U_{l}^{O}}\begin{pmatrix}W_{l}^{O,A}&0\\ 0&W_{l}^{O,B}\end{pmatrix}{\color[rgb]{1,0,0}M_{l}^{O}}\,.$
This process results in $Y_{l}^{*}\in\mathbb{R}^{T\times d}$, serving as the input for the subsequent MLP block at layer $l$.
B.1.3 The Multi-Layer Perceptron block
The Foldable SuperNet of the $l$ MLP block receives $Y_{l}^{*}\in\mathbb{R}^{T\times d}$ as input, which are the merged features of the previous attention block. It learns the layer norm parameters of the MLP block ${\color[rgb]{1,0,0}\alpha^{*}_{l},\theta^{*}_{l}}\in\mathbb{R}^{d}$, which, as usual, acts on the compressed dimension:
$\bar{Y}_{l}^{*}=\text{LN}_{{\color[rgb]{1,0,0}\alpha^{*}_{l},\theta^{*}_{l}}}(Y_{l}^{*})\,.$
After the layer norm, the ViT’s MLP block applies a sequence of operations: a linear layer, an activation function, and another linear layer. Our Foldable SuperNet mimics this process and uses $M$ and $U$ matrices to both compress and reconstruct the features at each stage, akin to the approach described in Section 2.1. After these operations, a skip connection is applied on the compressed dimension.
$X_{l+1}^{*}=Y_{l}^{*}+\sigma\left(\bar{Y}_{l}^{*}{\color[rgb]{1,0,0}U_{l}^{1}}\begin{pmatrix}W_{l}^{1,A}&0\\ 0&W_{l}^{1,B}\end{pmatrix}{\color[rgb]{1,0,0}M_{l}^{1}}\right){\color[rgb]{1,0,0}U_{l}^{2}}\begin{pmatrix}W_{l}^{2,A}&0\\ 0&W_{l}^{2,B}\end{pmatrix}{\color[rgb]{1,0,0}M_{l}^{2}}\,.$
Where ${\color[rgb]{1,0,0}U_{l}^{1},\,U_{l}^{2}}\in\mathbb{R}^{d\times 2\cdot d}$, ${\color[rgb]{1,0,0}M_{l}^{1},\,M_{l}^{2}}\in\mathbb{R}^{2\cdot d\times d}$. $X_{l+1}^{*}\in\mathbb{R}^{T\times d}$ then serves as the input for the $l+1$ attention block.
B.2 Training the Foldable SuperNet
In the case of merging two ViTs $A$ and $B$, $D^{A}$ and $D^{B}$ are defined as small subsets of training data from tasks $A$ and task $B$ respectively.
Our objective is to define a global optimization problem for training the Foldable SuperNet. As in the case of linear layers, we aim to reconstruct the features from the last representation layer (just before the classification head) of the original ViTs. $I_{\text{img}}^{k}$ represents an input image from task $k$, and $f^{k}_{L}(I_{\text{img}})\in\mathbb{R}^{d}$ represents the features from the last representation layer of the original model fine-tuned on task $k$, created from the input $I_{\text{img}}$. Observe that $f^{k}_{L}$ is the CLS token after being processed by the transformer and various post-processing stages that should also be merged (for instance, final layer normalization and a linear projection layer).
We will define the output of the Foldable SuperNet as $\tilde{f}_{L}(I_{\text{img}})\in\mathbb{R}^{2\cdot d}$, which is a reconstruction attempt for $f^{A}_{L}(I_{\text{img}})||f^{B}_{L}(I_{\text{img}})\in\mathbb{R}^{2\cdot d}$. Also, $\tilde{f}_{L}(I_{\text{img}})[k]$ will note the reconstruction attempt for model $k$ features. Then the loss function will be:
$L_{\text{out}}=\sum_{k}\,\mathbb{E}_{I_{\text{img}}^{k}\sim D^{k}}\,\left\|f_{L}^{k}(I_{\text{img}}^{k})-\tilde{f}_{L}(I_{\text{img}}^{k})[k]\right\|_{2}^{2}\,.$
This implies that for the input $I_{\text{img}}^{k}$ belonging to task $k$, we will only learn from the features of the model trained on this task, and not for example from the features that the model $j$ created $F^{j}(I_{\text{img}}^{k})$. This loss differs from the one used in the MLP case, where each layer attempts to reconstruct the features that both models create from the input $I_{\text{img}}^{k}$, regardless of the task it belongs to. This method was adopted for both FS-Merge and KD when merging ViTs, as we found it performed better.
In the case of merging ViTs, this global approach worked much better than addressing a series of local problems for each block, as was done in the MLP case. For more details, please refer to Appendix H.1.
B.3 Folding the Foldable SuperNet
Our next step after learning involves folding the Foldable SuperNet. This procedure aims to create a merged ViT that operates within the same dimensionality as the original models. The layer norm parameters acquired through our optimization process will be directly used in the merged model, as they already work in the merged dimension.
The folding operation (Eq. 5 ) follows the methodology outlined in Section 2. For instance, within the pre-processing block, the new merged projection weights and positional embeddings will be defined as follows:
$W^{\text{in},*}=(W^{\text{in},A}||W^{\text{in},B})M_{\text{in}}\,,$
$\text{emb}^{*}=(\text{emb}^{A}||\text{emb}^{B})M_{\text{in}}\,.$
Also, taking the attention block at layer $l$ as an example, the merged query weights will be:
$W^{Q,*}=U_{l}\begin{pmatrix}W_{l}^{Q,A}&0\\ 0&W_{l}^{Q,B}\end{pmatrix}M_{l}^{Q}\,.$
The other weights will be folded in a similar manner. Intuitively, this folding operation creates a merged model that, “under the hood”, uses $U$ to reconstruct the original features from the previous layer, applies the original weights, and then uses $M$ to merge those features again, all with the same complexity as each of the original models.
B.4 Merge tasks sequentially with FS-Merge seq.
Using the global version of FS-Merge on large models like transformers comes with a significant resource cost. As we have shown, the number of learnable parameters is smaller than in the distillation case, due to our modeling of the $M$ and $U$ matrices as a concatenation of diagonal matrices plus a low-rank matrix (Eq. 12). However, we still need to retain the frozen original weights of all models in memory, leading to increased memory and compute resource demands when merging multiple models.
To address this issue, we introduce FS-Merge Seq., which merges the models sequentially. For example, if we wish to merge $n$ models, we start by merging the first two models using the global FS-Merge (Eq. 4), with the features of these two original models as targets. The $M$ and $U$ matrices are still modeled as a concatenation of diagonal matrices plus a low-rank matrix, and we apply the “First” initialization (initialized from the first model).
After merging the first two models, we use global FS-Merge again to merge the resulting model (capable of solving the first two tasks) with the third model. In this phase, we use the features of all models seen so far (the first, second, and third models) as targets, and initialize from the merged model obtained in the previous step. This process is repeated, merging the previous merged model with a new original model at each step, until all $n$ models are merged. Note that In FS-Merge Seq., at each phase, we only merge and load the weights of two models (the previous merged model and a new original model), even though all features and models are utilized.
Experiments merging groups of four and five ViTs demonstrate that FS-Merge Seq. requires significantly less memory and compute resources, and merges models faster compared to regular FS-Merge. While this approach results in slightly lower accuracy compared to regular FS-Merge, FS-Merge Seq. still achieves better performance than distillation in most cases.
Appendix C Time and memory complexity analysis
For memory complexity analysis, let’s consider the merging of $n$ fully connected layers with weights $W\in\mathbb{R}^{d\times d}$ to one layer. When using distillation to merge these layers, $d^{2}$ learnable parameters are required as we train a single weight matrix for all $n$ models.
In the FS-Merge case, we will have $2nd^{2}$ learnable parameters in the $M$ and $U$ matrices. Additionally, we must hold $nd^{2}$ frozen weights in memory, which comes with reduced cost compared to learnable parameters (as we do not need to compute gradients for these matrices, and the optimizer does not need to save their moments). Nevertheless, this ‘vanilla’ version of FS-Merge is much more memory-intensive compared to Distillation.
To mitigate this issue, we suggest two additional versions of FS-Merge. In the first one, we parameterize the $M$ and $U$ matrices as a concatenation of diagonal matrices plus a low-rank matrix with a rank of $r$ (Eq. 12), resulting in $2(nrd+nd+rd)$ learnable parameters. As demonstrated in the transformer case, small values of $r$ are sufficient to outperform distillation, which also results in fewer learnable parameters compared to distillation. However, we still need to retain $nd^{2}$ frozen weights in memory.
Our final and most efficient version is FS-Merge seq. (Appendix B.4), which merges a pair of models at each stage. Thus, we use the previous calculation with $n=2$, which results in $6rd+4d$ learnable parameters and $2d^{2}$ frozen weights in memory at each stage. FS-Merge seq. comes with a small cost to performance, but still outperform distillation (and see Section 3.2, Appendix F.2).
Table 6 presents the total time and the number of optimized parameters when merging a group of four ViT-B-16 models with 100 original images and 1000 augmented images from each dataset. FS-Merge seq. used with a low rank of 16. The per-task accuracy is also shown.
Table 6: Measuring the total time and the number of optimized parameters, while merging a group of four ViT-B-16 with 100 original images and 1000 augmented images from each dataset. The per-task test accuracy is also reported. The merged models are the models fine-tuned on RESISC45, EuroSAT, CIFAR10, and MNIST.
Method Merging time #Parameters Optimized Accuracy
Average $\sim$ 4 Seconds 0 8.33
SLERP $\sim$ 4 Seconds 0 8.69
RegMean $\sim$ 3 Minutes 0 8.33
Opt $\sim$ 18 Minutes 0 8.76
Distillation $\sim$ 1.9 Hours  [^1] M 86.86
FS-Merge diagonal $\sim$ 3.2 Hours 900K 87.35
FS-Merge low rank $\sim$ 3.6 Hours 60M 91.54
FS-Merge seq. $\sim$ 2.2 Hours 18M 90.94
Appendix D Pre-training Vision Transformers
D.1 Training Vision Transformers from scratch
We pre-trained ViT-B-16 and ViT-L-14 (dosovitskiy2021an; steiner2022how; touvron2021training) on ImageNet-1K (deng2009imagenet). These models were initialized from distinct random seeds and exposed to training data in different orders. Following a setting similar to other merging works (ilharco2023editing; stoica2024zipit), a frozen classification head derived from CLIP’s (radford2021learning) label embeddings was used, in order to make the outputs space of the ViTs similar. Training and merging ViTs with learned classification heads are left for future research. It is important to mention that the classification heads are not being merged.
We pre-trained the ViTs following common practices (touvron2021training), such as augmentations, MixUp (zhang2018mixup), distillation using resnet152 (he2016deep), cross-entropy loss, and a cosine scheduler with a single cycle and a warmup.
D.2 Datasets and fine-tuned Vision Transformers
After obtaining different pre-trained ViTs, we fine-tuned each ViT on a different downstream task. Following the approach outlined in (ilharco2023editing), we fine-tuned on a range of datasets including Cars (krause20133d), DTD (cimpoi2014describing), EuroSAT (helber2019eurosat), GTSRB (stallkamp2011german), MNIST (lecun1998mnist), RESISC45 (cheng2017remote), SVHN (netzer2011reading), CIFAR10, and CIFAR100 (krizhevsky2009learning). We used the existing datasets’ training set, validation set, and test set. If there wasn’t a validation set, one was created by using 15% of the training set.
All models were fine-tuned with a batch size of 256, a learning rate of $1e^{-5}$, cross-entropy loss, and a cosine scheduler using a single cycle with a warm-up phase. As done in previous works (ilharco2023editing; stoica2024zipit), the classification head was frozen and obtained from CLIP embeddings (radford2021learning) of the label names. In Table 7, we present the dataset details and the test accuracy of the fine-tuned models on their respective tasks. The models’ accuracies on tasks they were not fine-tuned for are as good as a random guess.
Table 7: Dataset details and the test accuracy of the fine-tuned ViT-B-16 and ViT-L-14.
Dataset Train size Val size Test size #Classes ViT-B-16 test acc. ViT-L-14 test acc.
EuroSAT 18,700 3,300 5,00 10 99.06 98.10
GTSRB 22,644 3,994 12,630 43 98.32 97.88
Cars 6,923 1,221 8,041 196 86.12 93.32
CIFAR-10 42,500 7,500 10,000 10 97.33 98.9
CIFAR-100 42,500 7,500 10,000 100 85.61 92.74
DTD 1,880 1,880 1,880 47 64.04 64.57
MNIST 51,000 9,000 10,000 10 99.66 99.68
RESISC45 18,900 6,300 6,300 45 93.46 93.95
SVHN 62,269 10,988 26,032 10 96.78 96.63
Appendix E Merging experiments details and hyperparameters
E.1 Merging Multi-Layer Perceptrons
We divided the MNIST dataset (lecun1998mnist) into two subsets: the first containing labels 0 to 4 and the second containing labels 5 to 9. Each subset was further split into training, validation, and testing sets, with the validation set comprising 10% of the original training set. A similar approach was adopted for the Fashion-MNIST dataset (xiao2017fashion).
We trained MLPs on each of the MNIST subsets, utilizing SGD with a learning rate of 0.01, a batch size of 10, and cross-entropy loss. The training duration was set to 1 epoch. However, if the number of hidden layers exceeded four, or the hidden dimension surpassed 256, the training was extended to 2 epochs. For the Fashion MNIST dataset (xiao2017fashion), the same hyperparameters were used, with adjustments made only to the number of epochs, increased to 10, and the learning rate, reduced to 0.001.
Data. All the merging methods, with the exception of simple weight averaging, need features generated from the MLPs. Those were created by using unlabeled data from the training sets. 64 images from each split dataset were utilized, resulting in total of 128 images. Additionally, we normalized the target features of the two models to the same scale in FS-Merge and in distillation, because it led to an improvement in accuracy. Note that this will not hurt the performance of the merged model during inference, as the scale of the last layer’s features does not change the prediction.
Hyperparameters for the merging methods. The hyperparameters for the merging methods were determined separately for MNIST and Fashion MNIST. We chose the hyperparameters that maximize the per-task accuracy on the validation set when merging two MLPs with two hidden layers and a hidden width of 128. We then used those hyperparameters for merging MLPs with different depths and widths.
We will now outline the hyperparameters grid used for the hyperparameter search in each merging method. We used GD optimizer in FS-Merge and Distillation (so the batch size is 128, the whole data). The step learning rate scheduler employs two learning rate drops with $\gamma=0.9$, whereas the Cosine scheduler utilizes a single cycle with a warmup length of 20 epochs.
FS-Merge.
  • initialization type: “random”
  • num epochs: [1k, 5k, 10k, 15k (MNIST), 20k (Fashion MNIST), 25k]
  • learning rate: [0.3, 0.1, 0.03, 0.01, 0.003]
  • momentum: [0.9, 0.8]
  • scheduler: [step lr, cosine]
FS-Merge global.
  • initialization type: [“First”, “Average” (the average of the original models), “random”]
  • num epochs: [200, 400, 1k (MNIST), 1.5k (Fashion MNIST), 5k, 10k, 15k]
  • learning rate: [0.3, 0.1, 0.03, 0.01, 0.003]
  • momentum: [0.9, 0.8]
  • scheduler: [step lr, cosine]
FS-Merge global, ZipIt init.
  • initialization type: [“ZipIt”]
  • num epochs: [200, 400, 1k, 1.5k, 5k, 10k, 15k]
  • learning rate: [0.3, 0.1, 0.03, 0.01, 0.003]
  • momentum: [0.9, 0.8]
  • scheduler: [step lr, cosine]
Distillation.
  • initialization type: [“First”, “Average” (the average of the original models), “random”]
  • num epochs: [200, 400, 1k (MNIST), 1.5k (Fashion MNIST), 2k, 5k]
  • learning rate: [0.3, 0.1, 0.03, 0.01, 0.003]
  • momentum: [0.9, 0.8]
  • scheduler: [step lr, cosine]
RegMean.
  • $\alpha$: [1.0, 0.9 (MNIST), 0.8 (Fashion MNIST), 0.7, 0.6, 0.5, 0.4, 0.3, 0.2, 0.1]
E.2 Merging Vision Transformers
Data and augmentations. We took a set of images from the training set and expanded it using augmentations. Specifically, these augmentations were used: Random Crop; Random Horizontal Flip; Random choice between grayscale, Solarization, and GaussianBlur; and MixUp (zhang2018mixup). Then this dataset is used to create features, where features of the model fine-tuned on task $A$ were generated only from images of this task (and their augmentations). For efficiency and reproducibility, we saved these features and then used them for all our merging methods. In the case of FS-Merge and distillation, a single epoch means using the entire features dataset once, including the features created by augmentations.
Hyperparameter. Three types of hyperparameter experiments were conducted: merging pairs of ViT-B-16 in a low-data scenario, merging groups of four ViT-B-16, and merging pairs of ViT-L-14. For each experimental setting, a specific group of fine-tuned models was selected and a hyperparameter search was performed. The hyperparameters that maximized the per-task validation accuracy for this group were chosen and then applied when merging other model groups in this setting. The settings can be seen in Table 8.
Table 8: The different experimental settings. “#Original Images” refers to the number of original images taken from the training set per dataset. “#Augmented Images” refers to the number of augmented images created per dataset. The fine-tuned models refer to the models used for the hyperparameter search. When C10 = CIFAR10, G = GTSRB, R = RESISC45, S = SVHN.
Setting What is merged #Original Images #Augmented Images Total #Images fine-tuned models
a 2 ViT-B-16 16 800 1,632 R, C10
b 4 ViT-B-16 100 1000 4,400 R, C10, S, G
c 2 ViT-L-14 100 1000 2,200 R, C10
For all methods that require training (FS-Merge and distillation), a batch size of 128 was used, along with a cosine scheduler that utilized a single cycle with a warmup phase, an ADAMW optimizer with a weight decay of 0.001, and initialization from the first model (“First”). Similar to the MLP experiments, for the distillation baseline, the target features were scaled to an L2 norm of 0.5. However, in this specific setting, this scaling proved unhelpful for FS-Merge and was therefore not utilized. The hyperparameter grid used for the hyperparameter search will now be outlined.
FS-Merge, concatenation of diagonal matrices, without a low rank matrix.
  • epochs: [30, 100 (c), 200 (a, b), 300, 400]
  • lr: [0.1, 0.01 (c), 0.001 (a, b), 0.0001]
FS-Merge, concatenation of diagonal matrices + low rank matrix.
  • epochs: [30, 100, 200, 300, 400]
  • lr: [0.1, 0.01, 0.001, 0.0001, 0.00001]
FS-Merge seq. (Appendix B.4). “epochs” refers to the number of epochs used for all iterations except the last one. “last iteration epochs” denotes the number of epochs applied in the last iteration, which involves merging the final model with the model obtained from the previous iteration.
  • epochs: [10, 50, 100]
  • last iteration epochs: [50, 100, 200, 300]
  • lr: [0.0001]
Distillation.
  • epochs: [30, 100 (a, c), 200, 300 (b), 400]
  • lr: [0.1, 0.01, 0.001, 0.0001, 0.00001]
RegMean (jin2023dataless). $\alpha$ is a factor which decrease the non-diagonal items of the inner product matrices in the RegMean solution.
  • $\alpha$: [1.0, 0.9 (a, c), 0.8, 0.7 (b), 0.6, 0.5, 0.4, 0.3, 0.2, 0.1]
“Opt” (imfeld2023transformer). The hyperparameters include the filter type (which determines the feature tokens used by the optimal transport solver) and $\lambda$, a regularization term for the solver (Sinkhorn-Knapp algorithm). Lower values of $\lambda$ result in harder alignment.
  • filter: [Only CLS, Full (c), Window 2, Window 4, Window 6, Window 8 (a, b), Window 10, Window 14]
  • $\lambda$: [0, 0.08 (a,c), 0.2 (b), 0.5]
Appendix F Additional results
F.1 Merging Multi-Layer Perceptrons
Additional results were obtained by merging pairs of MLPs, using 64 images from each task to create features. Each experiment was replicated five times with different random seeds. Note that we do not merge the last linear layer (the classification head).
FS-Merge. Our method was tested in five variants. FS-Merge is the local version (Eq. 3), which trains a Foldable SuperNet for each layer $l$ independently, using the original models’ pre-activation features as inputs $z_{l}^{A},z_{l}^{B}$. FS-Merge global is the global version of our method (Eq. 4), training a Foldable SuperNet for all the layers together to reconstruct the features of the final representation layer $f_{L}^{A},f_{L}^{B}$. For both of these versions, we also tested a variant where we initialized the Foldable SuperNet’s $M$ and $U$ with the solutions of ZipIt (stoica2024zipit), and then continued to optimize them (FS-Merge ZipIt init and FS-Merge global ZipIt init). FS-Merge no cross compresses each of the two MLPs ($A$ and $B$) individually to half of their widths using a local FS-Merge and then concatenates those two compressed models. This means that neurons between these two models cannot be merged.
We also tried a local FS-Merge version where the $l$-th Foldable SuperNet layer uses the reconstructed features from the previous layer $\tilde{z}_{l}^{A},\tilde{z}_{l}^{B}$ as inputs, but it achieved the same accuracy as the regular local FS-Merge.
Table 9 presents the per-task accuracy for merging MLPs trained on half of the MNIST dataset (lecun1998mnist), with varying hidden width. Table 10 and Table 11 present the joint accuracy for fusing MLPs trained on half of the MNIST dataset, with variations in depth or hidden width, respectively. Similarly, Table 12 and Table 13 present the per-task accuracy for merging MLPs trained on half of the Fashion MNIST dataset (xiao2017fashion), again varying by depth or hidden width. The joint accuracy for these models are detailed in Table 14.
As shown in the experiments, our method, especially when using ZipIt initialization, demonstrates SOTA results across all settings and outperforms ensembles in some cases. It also appears that FS-Merge achieves better per-task accuracy, while global FS-Merge achieves better joint accuracy. Furthermore, ZipIt stands out as a strong baseline.
Table 9: Merging pairs of MLPs, each initialized differently and trained on distinct halves of the MNIST dataset. These MLPs have a single hidden layer, with the Hidden width varying from 16 to 1024. Each experiment was replicated five times. We present the average per task accuracy on the test set, along with the standard deviation.
Merge Method Hidden width
16 64 128 512 1024
Original Models 96.34 $\pm$ 0.3 96.77 $\pm$ 0.1 96.8 $\pm$ 0.1 97.8 $\pm$ 0.1 98.0 $\pm$ 0.1
Ensemble 84.47 $\pm$ 4.3 93.19 $\pm$ 1.8 94.0 $\pm$ 0.6 96.4 $\pm$ 0.3 96.5 $\pm$ 0.3
average 83.88 $\pm$ 3.9 93.05 $\pm$ 1.7 94.1 $\pm$ 0.4 96.6 $\pm$ 0.1 96.8 $\pm$ 0.1
RegMean 88.12 $\pm$ 3.6 95.06 $\pm$ 1.2 95.2 $\pm$ 0.2 96.7 $\pm$ 0.1 96.9 $\pm$ 0.2
ZipIt 91.84 $\pm$ 1.9 96.06 $\pm$ 0.12 96.3 $\pm$ 0.2 97.6 $\pm$ 0.1 97.7 $\pm$ 0.1
Distillation 82.69 $\pm$ 4.7 91.41 $\pm$ 2.5 93.0 $\pm$ 0.7 95.7 $\pm$ 0.4 96.0 $\pm$ 0.3
FS-M 76.7 $\pm$ 29.0 95.87 $\pm$ 0.1 95.8 $\pm$ 0.2 96.1 $\pm$ 0.2 95.9 $\pm$ 0.2
FS-M, ZipIt init 95.32 $\pm$ 0.4 96.50 $\pm$ 0.1 96.6 $\pm$ 0.1 97.5 $\pm$ 0.1 97.8 $\pm$ 0.1
FS-M no cross 63.8 $\pm$ 11.2 96.18 $\pm$ 0.1 96.1 $\pm$ 0.1 96.6 $\pm$ 0.1 96.4 $\pm$ 0.1
FS-M global 11.67 $\pm$ 4.2 95.66 $\pm$ 0.3 95.7 $\pm$ 0.1 96.3 $\pm$ 0.2 96.4 $\pm$ 0.2
FS-M global, ZipIt init 11.84 $\pm$ 4.6 96.44 $\pm$ 0.1 96.6 $\pm$ 0.1 97.6 $\pm$ 0.1 97.8 $\pm$ 0.1
Table 10: Merging pairs of MLPs, each initialized differently and trained on distinct halves of the MNIST dataset. These MLPs contain 128 neurons per hidden layer, with the number of hidden layers varying from 1 to 6. Each experiment was replicated five times. We present the average joint accuracy on the test set, along with the standard deviation.
Merge Method Number of Hidden Layers
1 2 3 4 6
Ensemble 80.20 $\pm$ 1.7 83.25 $\pm$ 3.3 84.6 $\pm$ 2.1 83.9 $\pm$ 1.9 83.1 $\pm$ 1.5
average 78.96 $\pm$ 0.7 69.83 $\pm$ 6.2 57.4 $\pm$ 5.2 38.2 $\pm$ 10.2 11.5 $\pm$ 2.9
RegMean 82.81 $\pm$ 1.5 80.38 $\pm$ 4.6 75.3 $\pm$ 2.8 74.7 $\pm$ 2.2 63.0 $\pm$ 7.0
ZipIt 85.33 $\pm$ 1.1 84.49 $\pm$ 1.9 81.5 $\pm$ 1.4 80.0 $\pm$ 1.4 80.2 $\pm$ 3.2
Distillation 77.32 $\pm$ 1.5 79.2 $\pm$ 3.96 79.6 $\pm$ 2.1 78.2 $\pm$ 3.6 74.4 $\pm$ 2.0
FS-M 84.44 $\pm$ 0.6 84.92 $\pm$ 1.0 82.7 $\pm$ 2.2 79.5 $\pm$ 3.1 79.1 $\pm$ 3.1
FS-M, ZipIt init 86.18 $\pm$ 0.6 86.57 $\pm$ 1.4 83.9 $\pm$ 3.1 79.7 $\pm$ 3.8 80.9 $\pm$ 3.6
FS-M no cross 85.39 $\pm$ 0.6 85.88 $\pm$ 1.3 86.0 $\pm$ 1.0 84.6 $\pm$ 1.1 83.8 $\pm$ 0.4
FS-M global 83.67 $\pm$ 0.7 84.12 $\pm$ 1.1 83.9 $\pm$ 1.0 82.8 $\pm$ 1.1 81.6 $\pm$ 0.7
FS-M global, ZipIt init 86.18 $\pm$ 0.7 86.40 $\pm$ 1.2 86.3 $\pm$ 1.2 84.7 $\pm$ 1.3 84.1 $\pm$ 0.9
Table 11: Merging pairs of MLPs, each initialized differently and trained on distinct halves of the MNIST dataset. These MLPs have a single hidden layer, with the hidden width varying from 16 to 1024. Each experiment was replicated five times. We present the average joint accuracy on the test set, along with the standard deviation.
Merge Method Hidden width
16 64 128 512 1024
Ensemble 58.86 $\pm$ 7.6 77.60 $\pm$ 5.5 82.1 $\pm$ 1.8 85.7 $\pm$ 0.5 87.0 $\pm$ 0.7
average 61.65 $\pm$ 8.0 77.16 $\pm$ 4.1 81.0 $\pm$ 2.2 84.4 $\pm$ 1.4 86.2 $\pm$ 0.6
RegMean 66.73 $\pm$ 7.7 81.54 $\pm$ 3.2 84.0 $\pm$ 1.6 85.2 $\pm$ 1.1 87.1 $\pm$ 1.0
ZipIt 66.89 $\pm$ 2.9 85.21 $\pm$ 1.0 83.9 $\pm$ 1.8 87.7 $\pm$ 1.5 88.0 $\pm$ 1.5
Distillation 57.15 $\pm$ 7.8 74.98 $\pm$ 4.9 79.6 $\pm$ 1.8 83.4 $\pm$ 0.6 85.0 $\pm$ 0.6
FS-M 59.1 $\pm$ 26.2 84.30 $\pm$ 1.0 83.9 $\pm$ 1.0 85.6 $\pm$ 0.9 85.5 $\pm$ 0.5
FS-M, ZipIt init 77.52 $\pm$ 4.2 85.91 $\pm$ 1.1 86.3 $\pm$ 0.7 88.5 $\pm$ 1.0 89.3 $\pm$ 0.5
FS-M no cross 54.00 $\pm$ 9.3 85.34 $\pm$ 1.5 85.0 $\pm$ 1.2 86.5 $\pm$ 0.8 86.5 $\pm$ 0.4
FS-M global 9.46 $\pm$ 0.1 83.72 $\pm$ 1.6 83.5 $\pm$ 1.1 85.0 $\pm$ 1.0 85.7 $\pm$ 0.5
FS-M global, ZipIt init 9.63 $\pm$ 0.1 85.99 $\pm$ 1.3 86.1 $\pm$ 0.9 88.6 $\pm$ 0.9 89.4 $\pm$ 0.4
Table 12: Merging pairs of MLPs, each initialized differently and trained on distinct halves of the Fashion MNIST dataset. These MLPs contain 128 neurons per hidden layer, with the hidden depth varying from 1 to 6. Each experiment was replicated five times. We present the average per task accuracy on the test set, along with the standard deviation.
Model Number of Hidden Layers
1 2 3 4 6
Original Models 90.45 $\pm$ 0.14 90.53 $\pm$ 0.1 90.4 $\pm$ 0.1 89.9 $\pm$ 0.1 83.0 $\pm$ 2.2
Ensemble 87.31 $\pm$ 1.93 88.68 $\pm$ 0.6 86.2 $\pm$ 1.9 86.9 $\pm$ 2.0 77.7 $\pm$ 4.0
average 86.04 $\pm$ 2.40 78.20 $\pm$ 5.1 58.9 $\pm$ 10.3 47.8 $\pm$ 5.4 22.4 $\pm$ 3.0
RegMean 89.29 $\pm$ 0.49 86.29 $\pm$ 0.2 81.3 $\pm$ 3.3 76.7 $\pm$ 4.4 64.1 $\pm$ 5.5
ZipIt 89.24 $\pm$ 0.44 87.40 $\pm$ 0.5 85.4 $\pm$ 2.9 83.2 $\pm$ 1.4 70.2 $\pm$ 8.5
Distillation 86.94 $\pm$ 1.49 87.84 $\pm$ 0.8 83.0 $\pm$ 1.6 83.2 $\pm$ 2.4 69.2 $\pm$ 3.9
FS-M 89.86 $\pm$ 0.13 89.80 $\pm$ 0.1 89.1 $\pm$ 0.2 88.4 $\pm$ 0.5 73.1 $\pm$ 4.1
FS-M, ZipIt init 90.20 $\pm$ 0.12 90.28 $\pm$ 0.1 90.0 $\pm$ 0.1 89.7 $\pm$ 0.2 80.0 $\pm$ 2.0
FS-M no cross 90.03 $\pm$ 0.13 90.20 $\pm$ 0.1 89.9 $\pm$ 0.1 89.3 $\pm$ 0.2 82.4 $\pm$ 2.2
FS-M global 89.85 $\pm$ 0.15 89.81 $\pm$ 0.2 89.3 $\pm$ 0.1 88.5 $\pm$ 0.2 78.5 $\pm$ 2.4
FS-M global, ZipIt init 90.04 $\pm$ 0.08 89.95 $\pm$ 0.1 89.5 $\pm$ 0.2 88.4 $\pm$ 0.2 62.5 $\pm$ 2.4
Table 13: Merging pairs of MLPs, each initialized differently and trained on distinct halves of the Fashion MNIST dataset. These MLPs have a single hidden layer, with the hidden width varying from 16 to 1024. Each experiment was replicated five times. We present the average per task accuracy on the test set, along with the standard deviation.
Merge Method hidden width
16 64 128 512 1024
Original Models 90.31 $\pm$ 0.06 90.43 $\pm$ 0.1 90.4 $\pm$ 0.1 90.6 $\pm$ 0.1 90.7 $\pm$ 0.1
Ensemble 72.71 $\pm$ 4.72 85.21 $\pm$ 3.3 86.9 $\pm$ 1.0 88.9 $\pm$ 0.4 87.0 $\pm$ 1.0
average 70.33 $\pm$ 6.56 83.90 $\pm$ 4.2 86.2 $\pm$ 1.4 88.5 $\pm$ 0.4 86.8 $\pm$ 0.8
RegMean 74.90 $\pm$ 8.31 87.33 $\pm$ 2.5 88.7 $\pm$ 1.5 89.1 $\pm$ 0.3 87.1 $\pm$ 1.3
ZipIt 77.39 $\pm$ 7.60 88.73 $\pm$ 0.7 89.4 $\pm$ 0.6 89.8 $\pm$ 0.2 88.7 $\pm$ 0.2
Distillation 74.06 $\pm$ 5.12 85.49 $\pm$ 2.9 86.7 $\pm$ 1.3 88.4 $\pm$ 0.6 86.5 $\pm$ 1.0
FS-M 89.06 $\pm$ 0.64 89.90 $\pm$ 0.1 89.8 $\pm$ 0.1 89.5 $\pm$ 0.1 89.1 $\pm$ 0.1
FS-M, ZipIt init 90.10 $\pm$ 0.19 90.25 $\pm$ 0.1 90.2 $\pm$ 0.1 90.3 $\pm$ 0.1 90.3 $\pm$ 0.1
FS-M no cross 49.2 $\pm$ 12.63 90.09 $\pm$ 0.1 90.1 $\pm$ 0.1 89.8 $\pm$ 0.1 89.5 $\pm$ 0.1
FS-M global 13.98 $\pm$ 4.88 89.87 $\pm$ 0.1 89.9 $\pm$ 0.1 89.6 $\pm$ 0.1 89.4 $\pm$ 0.1
FS-M global, ZipIt init 11.98 $\pm$ 3.97 89.98 $\pm$ 0.1 90.1 $\pm$ 0.1 90.2 $\pm$ 0.1 90.1 $\pm$ 0.1
Table 14: Merging pairs of MLPs, each initialized differently and trained on distinct halves of the Fashion MNIST dataset. These MLPs contain 128 neurons per hidden layer, with the hidden depth varying from 1 to 6. Each experiment was replicated five times. We present the average joint accuracy on the test set, along with the standard deviation.
Model Number of Hidden Layers
1 2 3 4 6
Ensemble 55.94 $\pm$ 4.14 56.28 $\pm$ 2.0 57.7 $\pm$ 2.9 56.9 $\pm$ 2.2 53.0 $\pm$ 0.9
average 54.71 $\pm$ 4.04 42.76 $\pm$ 6.4 33.6 $\pm$ 5.7 21.8 $\pm$ 1.8 8.9 $\pm$ 2.0
RegMean 56.94 $\pm$ 1.24 54.78 $\pm$ 2.1 54.7 $\pm$ 4.3 50.0 $\pm$ 4.5 51.6 $\pm$ 5.1
ZipIt 59.26 $\pm$ 1.66 60.01 $\pm$ 1.5 60.9 $\pm$ 2.4 59.6 $\pm$ 2.5 48.3 $\pm$ 7.3
Distillation 55.70 $\pm$ 2.49 54.33 $\pm$ 2.5 54.0 $\pm$ 3.9 52.6 $\pm$ 4.0 45.5 $\pm$ 3.9
FS-M 54.40 $\pm$ 0.29 49.91 $\pm$ 1.1 49.1 $\pm$ 1.5 47.3 $\pm$ 1.6 42.2 $\pm$ 2.3
FS-M, ZipIt init 54.13 $\pm$ 0.28 48.71 $\pm$ 1.0 48.1 $\pm$ 1.1 47.3 $\pm$ 1.2 44.3 $\pm$ 1.0
FS-M no cross 56.81 $\pm$ 0.23 56.19 $\pm$ 0.9 56.4 $\pm$ 1.3 56.7 $\pm$ 1.0 53.0 $\pm$ 0.8
FS-M global 56.63 $\pm$ 0.27 55.64 $\pm$ 0.9 55.8 $\pm$ 1.4 56.1 $\pm$ 0.9 51.9 $\pm$ 0.9
FS-M global, ZipIt init 60.46 $\pm$ 0.94 59.31 $\pm$ 0.3 60.1 $\pm$ 0.5 61.7 $\pm$ 2.3 40.5 $\pm$ 3.4
F.2 Merging Vision Transformers
For this section, C = Cars, C10 = CIFAR10, C100 = CIFAR100, D = DTD, E = EuroSAT, G = GTSRB, M = MNIST, R = RESISC45, S = SVHN.
Baselines. Our goal was to merge Vision Transformers (ViTs) that were initialized differently and trained on various tasks. We compared our method against “average” (wortsman2022model), a simple weight averaging technique; “SLERP” (shoemake1985animating), spherical linear interpolation; RegMean (jin2023dataless), which offers a closed-form solution by solving a linear regression problem for each linear layer; “Opt” (imfeld2023transformer), which uses optimal transport (knight2008sinkhorn) to align transformers, and can be viewed as a generalization of neuron alignment methods (ainsworth2023git) because it can find soft alignments as well as hard ones; and a distillation baseline (hinton2015distilling), which trains a single ViT to mimic the features of the last representation layer of the original models.
FS-Merge. We examined three versions of our method: FS-Merge low rank, where the $M$ and $U$ matrices were parameterized as a concatenation of diagonal matrices, plus a low rank matrix; FS-Merge diagonal, which is as “FS-Merge low rank” but with a low rank of 0; in the experiments of merging groups of four and five ViTs, we also used FS-Merge seq. (Appendix B.4), with a low rank of 16.
Additional results of the experiment involving the merging of pairs of ViT-B-16 can be seen in Table 15. This experiment examined our ability to merge models with a very low number of only 16 original images per dataset. An additional 800 images per dataset were created using augmentations. FS-merge low rank was used with a low rank of 12.
Table 16 presents experiments on merging groups of three ViT-B-16 models, each fine-tuned on different tasks, using 100 original images per dataset and 1000 augmented images per dataset. The FS-M low-rank method was applied with a rank of 28. All merging methods were used with the hyperparameters chosen for the experiment involving the merging of groups of four ViT-B-16 models.
Additional results of the experiment involving the merging of groups of four ViT-B-16 can be seen in Table 17 and Table 18. FS-merge low rank was used with a low rank of 32. In Table 19 we merged groups of five ViT-B-16, where FS-merge low rank was used with a low rank of 32. Note that solving five different tasks via merging is an extremely difficult challenge, and even ensemble struggles with it.
Table 20 and Table 21 present additional results from the experiment involving the merging of ViT-L-14 pairs. This experiment investigated how the merging methods scale to larger models, with FS-Merge Low Rank applied using a rank of 32.
Table 15: Merging pairs of ViT-B-16 with 16 original images from the training set and 800 augmented images from each dataset. We report the per-task and joint accuracy on the test set.
Merging Methods EuroSAT, CIFAR100 Cars, MNIST RESISC45, CIFAR10
Per-task Joint Per-task Joint Per-task Joint
Original models 92.33 - 92.89 - 95.39 -
Ensemble 90.25 83.72 90.44 89.63 92.17 86.06
Average 5.635 1.94 4.96 4.87 5.61 1.66
SLERP 5.88 2.87 7.12 4.34 4.82 1.76
RegMean 4.45 0.95 5.18 0.27 8.45 5.44
Opt 5.60 0.90 5.37 5.16 6.32 5.32
Distillation 72.02 66.34 80.14 63.28 65.47 59.64
FS-M diagonal 69.53 63.98 85.88 73.10 72.11 65.68
FS-M low rank 71.86 68.23 88.46 74.95 74.59 69.45
Table 16: Merging groups of 3 ViT-B-16 with 100 original images from the training set and 1000 augmented images from each dataset (resulting in a total of 300 original images and 3,000 augmented images). We report the per-task and joint accuracy on the test set. “Parameters Optimized” refers to the number of learnable parameters. We will denote: C = Cars, C10 = CIFAR10, C100 = CIFAR100, D = DTD, E = EuroSAT, G = GTSRB, M = MNIST, R = RESISC45, S = SVHN.
Merging Methods R, C100, S C100, C, M D, M, E
Per-task Joint Per-task Joint Per-task Joint #Parameters Optimized
Original models 91.95 - 90.46 - 87.58 - -
Ensemble 84.28 48.49 83.92 61.67 82.71 63.02 -
Average 4.02 0.34 3.77 0.25 7.76 3.39 0
SLERP 3.88 0.69 3.69 0.08 8.84 5.11 0
RegMean 3.81 0.83 4.76 0.20 7.17 0.71 0
Opt 3.95 0.52 4.01 0.13 7.22 2.36 0
Distillation 65.03 52.36 59.79 49.64 79.04 76.76  [^1] M
FS-M, diagonal 63.77 45.32 57.54 38.97 79.06 59.70 500K
FS-M, low rank 67.61 57.98 62.26 39.74 83.11 80.43 42M
Table 17: Merging groups of 4 ViT-B-16 with 100 original images from the training set and 800 augmented images from each dataset (resulting in a total of 400 original images and 3,200 augmented images). We report the per-task and joint accuracy on the test set. We will denote: C = Cars, C10 = CIFAR10, C100 = CIFAR100, D = DTD, E = EuroSAT, G = GTSRB, M = MNIST, R = RESISC45, S = SVHN.
Merging Methods G, M, C100, S E, G, C10, S R, E, C10, M
Per-task Joint Per-task Joint Per-task Joint
Original models 95.09 - 97.87 - 97.37 -
Ensemble 85.98 40.31 91.97 58.64 89.41 50.56
Average 5.12 0.05 8.57 1.44 8.33 0.68
SLERP 6.20 0.33 7.81 1.12 8.69 0.43
RegMean 6.32 0.25 11.29 1.54 8.33 1.23
Opt 5.17 0.16 11.53 3.90 8.76 1.62
Distillation 76.18 42.67 34.55 29.20 86.86 75.90
FS-Merge, diagonal 68.87 34.59 79.30 67.62 87.35 72.17
FS-Merge, low rank 72.93 39.48 85.69 82.10 91.54 77.63
FS-Merge seq. 74.11 39.63 84.64 75.93 90.66 79.68
Table 18: Merging groups of 4 ViT-B-16 with 100 original images from the training set and 800 augmented images from each dataset (resulting in a total of 400 original images and 3,200 augmented images). We report the per-task and joint accuracy on the test set. We will denote: C = Cars, C100 = CIFAR100, D = DTD, E = EuroSAT, G = GTSRB, M = MNIST, R = RESISC45, S = SVHN.
Merging Methods C, C100, R, E D, M, S, C100 E, D, R, M
Per-task Joint Per-task Joint Per-task Joint #Parameters Optimized
Original models 91.06 - 86.52 - 89.05 - -
Ensemble 77.56 64.78 73.72 29.98 78.58 42.25 -
Average 5.19 0.26 5.24 0.14 7.11 0.68 0
SLERP 3.15 0.19 5.37 0.25 7.47 0.59 0
RegMean 4.03 0.27 5.94 0.25 7.11 0.41 0
Opt 4.17 0.23 5.92 0.26 7.94 1.67 0
Distillation 59.28 53.04 66.22 36.50 53.52 45.85  [^1] M
FS-Merge, diagonal 60.35 46.83 60.69 29.01 56.51 46.76 600K
FS-Merge, low rank 69.45 60.48 67.44 36.81 67.78 59.80 60M
FS-Merge seq. 70.05 58.08 67.23 38.76 65.26 55.20 18M
Table 19: Merging groups of 5 ViT-B-16 with 100 original images from the training set and 800 augmented images from each dataset (resulting in a total of 500 original images and 4,000 augmented images). We report the per-task and joint accuracy on the test set. We will denote: C = Cars, C10 = CIFAR10, C100 = CIFAR100, D = DTD, E = EuroSAT, G = GTSRB, M = MNIST, R = RESISC45, S = SVHN.
Merging Methods R, M, D, S, C10 C, M, C100, E, R E, S, C, C10, G
Per-task Joint Per-task Joint Per-task Joint
Original models 90.25 - 92.78 - 95.52 -
Ensemble 75.06 35.98 76.77 50.06 82.69 65.32
Average 7.22 1.65 4.63 0.37 6.57 0.10
SLERP 7.40 0.40 6.33 0.62 6.34 0.59
RegMean 7.08 1.91 4.64 0.67 9.16 1.57
Opt 7.67 0.65 3.98 1.69 9.18 3.18
Distillation 74.87 63.99 65.72 56.22 77.04 64.27
FS-Merge, diagonal 73.30 59.82 67.16 50.49 70.27 59.00
FS-Merge, low rank 78.20 65.90 76.13 66.91 79.71 64.94
FS-Merge seq. 76.05 50.64 72.93 61.55 78.28 60.65
Table 20: Merging pairs of ViT-L-14 with 100 original images from the training set and 1000 augmented images from each dataset. We report the per-task and joint accuracy on the test set.
Merging Methods RESISC45, CIFAR10 GTSRB, RESISC45 Cars, EuroSAT
Per-task Joint Per-task Joint Per-task Joint
Original models 96.42 - 95.91 - 95.71 -
Ensemble 95.52 94.12 94.05 93.35 93.07 92.95
Average 11.21 3.83 1.92 1.10 9.36 1.97
SLERP 16.14 4.48 2.42 1.43 8.23 5.71
RegMean 13.14 8.61 4.56 2.69 6.91 6.04
Opt 10.89 3.06 3.60 2.46 3.07 0.28
Distillation 84.37 82.20 82.62 80.64 89.11 86.98
FS-M diagonal 82.93 77.91 76.94 73.83 91.14 82.54
FS-M low rank 85.82 83.11 80.39 78.43 93.22 90.04
Table 21: Merging pairs of ViT-L-14 with 100 original images from the training set and 1000 augmented images from each dataset. We report the per-task and joint accuracy on the test set.
Merging Methods RESISC45, SVHN DTD, GTSRB CIFAR100, EuroSAT
Per-task Joint Per-task Joint Per-task Joint
Original models 95.29 - 81.22 - 95.42 -
Ensemble 93.53 90.37 78.22 77.02 91.13 90.46
Average 8.32 1.07 2.66 1.09 9.87 0.82
SLERP 5.97 1.65 1.80 0.72 10.48 6.29
RegMean 12.95 3.32 4.83 2.98 13.14 8.61
Opt 5.14 4.26 4.03 2.95 3.54 0.52
Distillation 87.66 77.28 67.20 65.84 92.39 86.62
FS-Merge diagonal 85.93 77.27 63.79 61.36 91.99 84.8
FS-Merge low rank 88.10 79.33 69.23 67.22 92.41 89.74
F.3 Merging text Transformers
In this series of experiments, we aimed to evaluate FS-Merge on a different modality: merging differently initialized text transformers. Following (verma2024merging), we used five bert-base-uncased models from the MultiBERTs reproductions (devlin2018bert; sellam2022the), each trained from a different random initialization and with a different data ordering. We fine-tuned each model on a distinct classification task from the GLUE dataset (wang2018glue), using six tasks: RTE, QQP, MNLI, MRPC, SST-2, and QNLI. The same pre-trained BERT was fine-tuned on QQP and MRPC, so their merge was not evaluated. It is worth noting that ViT is a pre-LN Transformer, while BERT is a post-LN Transformer (xiong2020layer).
Baselines. We compared our method against two baselines: “average” (wortsman2022model), a simple weight averaging technique; and distillation (hinton2015distilling), which trains a single BERT to mimic the features of the last representation layer of the original models. We also reported the performance of the “original models”, representing the average accuracy of the models to be merged; and ensemble (ganaie2022ensemble), which averages the models’ outputs and then applies classification heads. Note that these last two are not valid merging methods as they use the original models directly.
Hyperparameters. We performed a hyperparameter search similar to the ViT case (Appendix E.2). We selected a pair of tasks, QQP and MNLI, and created a validation set for each task from the training set. We then fine-tuned a pair of differently initialized BERT models, one for each task. These two models were used to conduct a hyperparameter search for both distillation and FS-Merge using the same hyperparameter grid. The hyperparameters that maximized the average per-task validation accuracy for this pair were chosen and then applied when merging the other BERT pairs. For FS-Merge, the chosen hyperparameters were 400 epochs, learning rate of 0.0001, weight decay of 0, “first” initialization, and low rank of 12. For Distillation, the chosen hyperparameters were 300 epochs, learning rate of 0.0001, weight decay of 0, and “first” initialization. Both methods used batch size of 128.
200 data points were taken from each training set to create features for the merging methods. In Table 22, we report the per-task test accuracy of these experiments. We can see that, similarly to the ViT case, traditional merging methods like “average” fail on the challenging task of merging text transformers from different initializations. FS-Merge achieve SOTA results, outperforming the ensemble in some cases.
Table 22: Merging pairs of BERTs with 200 original texts from the training set of each dataset. We report the per-task accuracy on the test set.
Tasks Original models Ensemble Average Distillation FS-Merge
RTE, QQP 79.49 71.13 41.87 65.00 68.88
MNLI, MRPC 85.73 67.17 50.52 76.87 78.29
MRPC, QNLI 89.02 87.19 40.53 80.67 79.33
SST-2, RTE 80.15 71.64 51.06 75.45 76.00
MNLI, SST-2 88.39 84.78 43.75 82.81 83.24
RTE, QNLI 79.75 74.60 48.94 66.40 69.69
QNLI, QQP 91.01 71.63 56.32 82.40 83.03
SST-2, QNLI 91.68 90.44 49.38 82.92 83.90
MRPC, RTE 77.49 68.41 57.83 68.70 68.37
QNLI, MNLI 87.99 71.83 41.62 71.83 73.79
QQP, RTE 79.49 71.13 41.87 73.30 75.62
MRPC, MNLI 85.73 67.17 50.52 62.90 62.81
QQP, SST-2 91.41 86.59 55.48 87.49 87.94
MNLI, QQP 87.7 78.69 36.09 78.74 79.22
QNLI, MRPC 89.02 87.19 40.53 82.97 82.98
RTE, SST-2 80.15 71.64 51.06 69.20 73.39
QNLI, SST-2 91.68 90.44 49.38 87.27 88.09
RTE, MRPC 77.49 68.41 57.83 63.46 66.05
SST-2, MNLI 88.39 84.78 43.75 66.44 67.29
QQP, QNLI 91.01 71.63 563.2 82.16 82.79
F.4 Number of original training images - Vision Transformer
Here we present additional experiments which examine the effect of the number of original images versus augmented images on the ViT merged model. Original images, defined as those sourced from the training datasets of the models to be merged, were varied in number: 16, 64, 128, 256, 512, 1024. For each scenario, augmented images were generated to bring the total images per dataset to 1024, ensuring uniform dataset size across all cases. These images were then used to create features for the merging processes.
Ensemble, Distillation, diagonal FS-Merge (low rank of 0), and FS-Merge with low rank of 24 were applied to merge pairs of ViT-B-16 models. It should be noted that Ensemble is not considered a legitimate merging method. The per-task and joint accuracy on the test set were reported, employing the optimal hyperparameters identified in earlier chapters. The outcomes of this experiment are presented in Figure 5 and Figure 6. Note that in these experiments, different features were used compared to the old experiments, so the results may vary.
Figure 5: We used Ensemble, Distillation, and FS-Merge to merge pairs of models trained on EuroSAT and Cars (left), CIFAR100 and SVHN (center), RESISC45 and EuroSAT (right). We varied the number of original images per dataset and added augmentation images so the total number of images per dataset would be 1024. We present the per-task and joint accuracy.
Figure 6: We used Ensemble, Distillation, and FS-Merge to merge pairs of models trained on GTSRB and MNIST (left), DTD and SVHN (center), CIFAR100 and EuroSAT (right). We varied the number of original images per dataset and added augmentation images so the total number of images per dataset would be 1024. We present the per-task and joint accuracy.
F.5 Number of training data points - BERT
We examined the impact of the amount of data points taken on the merged BERT model. Training data points, sourced from the training datasets of the models to be merged, were adjusted to sizes of: 16, 64, 128, 256, 512, 1024. These samples were used to generate features for the merging process. Unlike the ViT case, we did not employ data augmentation, resulting in non-uniform dataset sizes in each case. To compensate, an additional 200 data points from each dataset were used to implement early stopping, preventing overtraining with smaller datasets.
We applied Ensemble, Distillation, and FS-Merge with a low rank of 12 to merge pairs of BERT base models. It is important to note that Ensemble is not considered a valid merging method. We reported per-task accuracy on the test set using optimal hyperparameters identified previously. The results of this experiment are shown in Figures 7.
Figure 7: Merging of BERT model pairs trained on RTE and SST-2 (left), QNLI and MRPC (center), MNLI and QQP (right) using Ensemble, Distillation, and FS-Merge. The graph shows per-task test accuracy across varying training data sizes. Three runs were conducted, each with a different random seed, to generate error bars and verify the statistical significance of the results.
Appendix G Ablation studies
G.1 FS-Merge ablation study
Table 23 shows an ablation study of merging a group of three ViT-B-16 using FS-Merge. This study involved three groups of models. We evaluated both the per-task average accuracy and joint accuracy on the test set.
We first used FS-Merge diagonal, where the $M,U$ were parameterized as a concatenation of diagonal matrices. We used random initialization, and created features using only the original 100 images from each dataset without augmentations (so a total number of 300 images). Then, 1000 more augmented images were added for each dataset. Then, “average initialization” for the $M,U$ was tested, meaning the Foldable SuperNet initialized the merged model from the average of the original ones; and also first initialization was tested, so the Foldable SuperNet initialized the merged model from the first original model. “Low rank” stand for FS-Merge with $M,U$ parametrized as a concatenation of diagonal matrices plus low-rank matrices (Eq. 12), with low rank of 24.
As can be observed, the “first” initialization leads to a better merged model compared to the average or random initialization. Additionally, this experiment shows the significance of creating more images through augmentations and using low rank in the $M,U$ matrices of the Foldable SuperNet.
Table 24 displays the per-task accuracy for each task of the merged model across RESISC45, CIFAR10, and EuroSAT. It can be seen that the “first” initialization improves not only the first task’s accuracy but also the accuracy across all tasks. This behavior recurs in every other group of models that we merged.
Our experiments suggest that in the ViT case, initialization is extremely important for training the Foldable SuperNet. When initialized randomly, FS-Merge does not converge, even with the addition of augmented data (and see Appendix H.2 discussing this issue). Only Foldable SuperNet initialized from the average or first model allows it to converge into a successful merged model.
Table 23: Ablation study comparing the effectiveness of different initialization and augmentation strategies on the merging of groups of three ViT-B-16. Only our method, FS-Merge, was used. We show the per-task and joint accuracy on the test set. “Aug” stands for Augmentations, and “Init” stands for Initialization. We will denote: C = Cars, C10 = CIFAR10, C100 = CIFAR100, D = DTD, E = EuroSAT, G = GTSRB, M = MNIST, R = RESISC45, S = SVHN.
Setting R, C10, E S, E, G C10, G, M
FS-Merge Init Aug Per-task Joint Per-task Joint Per-task Joint
Diagonal Random 7.43 0.77 11.05 1.18 7.71 1.18
Diagonal Average 25.87 10.15 21.42 12.19 14.71 4.25
Diagonal First 76.75 55.44 66.87 50.20 67.74 51.38
Low rank First 83.40 73.76 77.48 65.69 80.77 64.93
Diagonal First 86.04 73.47 82.15 71.18 76.45 57.13
Low rank First 89.17 79.34 85.96 75.38 86.41 66.84
Table 24: Merging three ViT-B-16 models, fine-tuned on RESISC45, CIFAR10 and EuroSAT, using diagonal FS-Merge. We are examining the effects of initialization and augmentation on the per-task accuracy of the test set.
FS-Merge details Original tasks
Initialization Augmentations Average Acc RESISC45 CIFAR10 EuroSAT
Average 25.87 3.73 22.94 50.94
First 76.75 89.4 56.24 84.62
Average 37.51 7.3 33.59 71.64
First 86.04 87.86 79.76 90.52
G.2 Distillation ablation study
Table 25 presents an ablation study on merging a group of three ViT-B-16 models using distillation. This study involved the same three groups of models used in the FS-Merge ablation study (Table 23). We followed Appendix G.1, and created features using 100 original images and 1000 augmented images from each dataset (resulting in a total of 3,300 images).
Table 25: Ablation study comparing the effectiveness of different initialization and augmentation strategies on the merging of groups of three ViT-B-16. Only distillation merge was used. We show the per-task and joint accuracy on the test set. We will denote: C = Cars, C10 = CIFAR10, C100 = CIFAR100, D = DTD, E = EuroSAT, G = GTSRB, M = MNIST, R = RESISC45, S = SVHN.
Distillation R, C10, E S, E, G C10, G, M
Initialization Augmentations Per-task Joint Per-task Joint Per-task Joint
Random 33.20 25.93 29.00 25.75 27.56 19.77
Random 39.64 30.83 34.39 30.13 33.17 24.69
Average 38.16 30.22 37.78 33.49 37.40 25.68
Average 41.78 32.76 55.35 46.20 52.84 39.54
RegMean 49.44 39.61 66.50 55.54 57.80 46.31
RegMean 58.53 47.84 77.18 66.42 72.55 57.42
First 81.56 71.32 77.24 65.02 78.38 67.76
First 84.14 73.98 84.46 73.19 84.30 70.92
We aimed to investigate how different initializations affect the performance of the distillation merge, also examining traditional merging baselines as initializations (such as average and RegMean). Additionally, we explored the impact of augmentations. We evaluated both the per-task average accuracy and the joint accuracy on the test set.
As observed, the “first” initialization leads to a superior merged model compared to all other initializations, including other merging baselines such as average and RegMean. Moreover, augmentations enhance performance in all cases.
In Table 26, we compare the best distillation and FS-Merge versions from the ablation studies (i.e., with “first” initialization), showing that FS-Merge outperforms distillation with and without augmentations.
Table 26: Comparing Distillation and FS-Merge (both with “first" init), with and without augmentations, while merging groups of three ViT-B-16. Features were created using 100 original images and 1,000 augmented images per dataset. We show the per-task and joint accuracy on the test set. We will denote: C = Cars, C10 = CIFAR10, C100 = CIFAR100, D = DTD, E = EuroSAT, G = GTSRB, M = MNIST, R = RESISC45, S = SVHN.
Setting R, C10, E S, E, G C10, G, M
Method Aug Per-task Joint Per-task Joint Per-task Joint
Distillation $\times$ 81.56 71.32 77.24 65.02 78.38 67.76
FS-Merge, Low rank $\times$ 83.40 73.76 77.48 65.69 80.77 64.93
Distillation $\surd$ 84.14 73.98 84.46 73.19 84.30 70.92
FS-Merge, Low rank $\surd$ 89.17 79.34 85.96 75.38 86.41 66.84
Appendix H Discussion
H.1 Merging Vision Transformers: Local VS. Global perspective
Observe that in the MLP merging method from Section 2.1, the local version of FS-Merge was used, which involves training a Foldable SuperNet individually for each layer. In the ViT case, local FS-Merge involves training each block of the Foldable SuperNet separately. For example, training a Foldable SuperNet that merges the first attention blocks, then training a Foldable SuperNet that merges the first MLP blocks, and so on. We found empirically that this approach leads to a poor solution in the ViT case, resulting in a dysfunctional merged model with accuracy nearly as poor as a random guess. Instead, we found that the global version of FS-Merge is much more effective in this case, involving training the entire Foldable SuperNet of the ViT to reconstruct the features of the last representation layer of the original models, $f_{L}$.
A few explanations exist for this issue. First, training the Foldable SuperNet in a local manner for the ViT, meaning block-wise, must be performed on very unnatural blocks, which “break” the transformer blocks. This is necessary because $M$ and $U$ matrices must be placed before or after a linear layer to allow them to be folded after training. Moreover, the attention score computation, layer normalization, and skip connections must be performed on the merged features (with the lower dimensionality). These conditions forced the “breaking” of existing ViT blocks, and, for example, required teaching the Foldable SuperNet of the attention block to reconstruct features that are within the next MLP blocks. This complicated structure probably have hindered the optimization process.
Second, the Foldable SuperNet consistently uses the merged features in the skip connection. This means that when learning block-wise, the features forwarded via skip connection to the next block are dramatically changed. These new merged features are very different from the original ones, which likely severely affected the optimization of the next Foldable SuperNet block.
H.2 Merging Vision Transformer with randomly initialized Foldable SuperNet
In our experiments, we found that smart initialization of the Foldable SuperNet is crucial for FS-Merge. As common in NNs, we first tried to initialize the $M,U$ matrices using a random Gaussian distribution dependent on the hidden dimension (he2015delving).
In the MLP case, a random initialization can work, but better results are achieved when using ZipIt as the initialization method for the Foldable SuperNet (and see Section 3.1). In the ViT case, We tested multiple scales for the random initialization, but could not find a setup that allowed the FS-Merge to converge into a functional merged model.
This led us to study smarter initializations, such as “average”. When merging $K$ models, a Foldable SuperNet that creates an average merge of the weights is achieved by setting all the $M,U$ matrices as follows:
$M=\begin{pmatrix}\frac{I}{K}\\ ...\\ \frac{I}{K}\\ \end{pmatrix},\ U=\begin{pmatrix}I&...&I\end{pmatrix}\,.$
When $I$ is the identity matrix. In the case of ViTs, the “first” initialization proved to be the most effective, involving initializing the Foldable SuperNet so it exclusively selects the weights of the first model. It can be achieved by setting all the $M,U$ matrices as follows:
$M=\begin{pmatrix}I\\ 0\\ ...\\ 0\\ \end{pmatrix},\ U=\begin{pmatrix}I&...&I\end{pmatrix}\,.$
By the end of the training, the “first” initialization results in a merged model with improved accuracy across all tasks, not just the task of the first model (see Appendix G.1 for more details). Surprisingly, averaging initialization also performed well, despite the fact that the average merge is not an effective merging method when combining ViTs trained from different initializations.
This effectiveness is the reason the Foldable SuperNet’s $M,U$ matrices were modeled as a sum of low-rank matrices plus a concatenation of diagonal matrices (rather than just a low-rank matrix as in LoRa (hu2022lora)). The concatenation of diagonal matrices enables the initialization of the $M,U$ matrices using those successful methods (“first”, “average”).
H.3 Using inner features when merging Vision Transformers
In line with several distillation studies (wu2021one; zagoruyko2017paying; heo2019knowledge; heo2019comprehensive; park2019feed; liu2020adaptive), we tried to use the inner features of the models to be merged as a regularization for FS-Merge and for distillation. We focused on the features obtained after the MLP block or after the attention block. The attention features in the $l$ block of model $k$, created from the input $I_{\text{img}}^{k}$, can be written as $f^{k}_{l,\text{att}}(I_{\text{img}}^{k})\in\mathbb{R}^{T\times d}$. Then, the “inner loss” for the global FS-Merge, when handling two tasks $A$ and $B$, can be defined as follows:
$L=L_{\text{out}}+\,\lambda\sum_{l\in C}\,\mathbb{E}_{I_{\text{img}}^{A}\sim D^{A}}\,\left\|f^{A}_{l,\text{att}}(I_{\text{img}}^{A})-\tilde{f}_{l,\text{att}}(I_{\text{img}}^{A})[A]\right\|_{2}^{2}+$
$\mathbb{E}_{I_{\text{img}}^{B}\sim D^{B}}\,\left\|f^{B}_{l,\text{att}}(I_{\text{img}}^{B})-\tilde{f}_{l,\text{att}}(I_{\text{img}}^{B})[B]\right\|_{2}^{2}\,.$
Table 27: The effect of inner loss on distillation, FS-Merge low rank, and FS-Merge full rank is demonstrated. Pairs of ViT-B-16, fine-tuned on RESISC45 and CIFAR10, were merged using 100 original images and 1000 augmented images per task. The per-task accuracy on the test set is presented.
Method Inner loss details Per-task Accuracy
Distillation $\lambda=0$ 86.41
Distillation $\lambda=0.1,n=\{5\}$ 83.22
FS-Merge, rank 12 $\lambda=0$ 89.78
FS-Merge, rank 12 $\lambda=0.1,n=\{5\}$ 86.39
FS-Merge, full rank No regularization 40.20
FS-Merge, full rank $\lambda=0.1,n=\{5\}$ 52.77
FS-Merge, full rank $\lambda=0.1,n=\{3,5,7,9\}$ 60.41
FS-Merge, full rank $\lambda=0.5,n=\{3,5,7,9\}$ 70.52
FS-Merge, full rank $\lambda=1,n=\{3,5,7,9\}$ 70.11
Where $L_{\text{out}}$ is the regular global reconstruction loss, which attempts to reconstruct the features of the two original models from the layer preceding the classification head (Appendix B.2). $\tilde{f}_{l,\text{att}}[k]$ are the reconstructed attention features of our Foldable SuperNet in block $l$ for model $k$. $C$ is the set of blocks from which we decided to extract features, and $\lambda$ is the regularization coefficient. $D^{A}$ and $D^{B}$ are defined as small subsets of data from the training data of tasks $A$ and task $B$ respectively.
This can easily be rewritten for distillation, only that for each $k$ we compare $f_{l,\text{att}}^{k}$ with $\tilde{f}_{l,\text{att}}$, which are the inner features of the student model in the matching block. Note that in the distillation method, this regularization forces the inner features $\tilde{f}_{l,\text{att}}$ to resemble the inner features of all the models to be merged (in our example, both $A$ and $B$). In contrast, in FS-Merge, after applying the $U$ matrix, the reconstruction of features from both models $A$ and $B$ are obtained, and each set of these features will be compared with its corresponding ground truth.
We tried multiple $\lambda$ values, and various $C$ sets, for features from the MLP or attention blocks. Yet, it seems to only detriment the performance of FS-Merge and distillation. Table 27 shows some of those experiments, merging pairs of ViT-B-16, fine-tuned on RESISC45 and CIFAR10, using 100 original images and 1000 augmented images per task.
It should be observed that specifically in the case of FS-Merge using full rank $M$ and $U$ matrices, the inner loss seems to improve the results. We conclude this because, in the full $M,U$ case, there is a very large number of learnable parameters, so a stronger regularization is needed. However, using full rank $M,U$ matrices in the ViT case is not recommended due to the very high memory and time complexity, and even with this regularization, it underperforms compared to Low rank FS-Merge.